### v1

# A Note on Restricted Selection Set from Random Interval

Abstract.
We study restricted selection sets of random intervals in $\mathbb{R}^{1}$ defined on a non-atomic probability space. Given a random interval $Y=[y_{L},y_{U}]$ and scalar constraints on the expectation or the median of admissible selections, we define the restricted selection set and establish its existence, basic structure and influence on bounding moments and quantiles. In particular, we give conditions under which any mean (or quantile) in the Aumann expectation range can be attained by a measurable selection. We characterize the induced ranges of means, medians, and event probabilities. The analysis is carried out in a minimal one-dimensional random-set framework inspired by the classical theory of Aumann integrals. We also outline extensions to higher-order moment and general quantile restrictions.
Key words and phrases:
Random sets, random intervals, Aumann expectation, selection, partial identification
1. Introduction
Let $(\Omega,\mathcal{A},\mathbf{P})$ be a complete non-atomic probability space and let $Y=[y_{L},y_{U}]$ be an interval-valued random closed set on $\mathbb{R}$, that is, a measurable map $Y:\Omega\to\mathcal{K}(\mathbb{R})$ taking values in compact intervals
$Y(\omega)=[y_{L}(\omega),y_{U}(\omega)],\qquad\omega\in\Omega,$
for some real-valued random variables $y_{L},y_{U}$ with $y_{L}\leq y_{U}$ almost surely. A measurable selection of $Y$ is a random variable $y:\Omega\to\mathbb{R}$ such that $y(\omega)\in Y(\omega)$ for $\mathbf{P}$–almost every $\omega$. The Aumann (selection) expectation of $Y$ is the closed set
$\mathbb{E}[Y]:=\overline{\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}^{1}(Y)\}},$ (1.1)
where $\operatorname{\mathbf{Sel}}^{1}(Y)$ denotes the family of integrable selections of $Y$; see [Aumann1965, Artstein1974, Molchanov2005]. In the one-dimensional interval case this set has the simple representation
$\mathbb{E}[Y]=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big],$ (1.2)
so that the possible expectations of selections $y\in\operatorname{\mathbf{Sel}}(Y)$ form a compact interval determined solely by the expectations of the endpoints.
In this note, we ask how the bounds in (1.2) change once one restricts attention to a non-empty subset of selections $\mathcal{S}\subset\operatorname{\mathbf{Sel}}^{1}(Y)$. Specifically, we consider subsets of $\operatorname{\mathbf{Sel}}^{1}(Y)$ satisfying elementary scalar constraints. Concretely, we consider restricted selection sets defined by moment or quantile conditions, such as
$\operatorname{\mathbf{Sel}}(Y\mid\kappa):=\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}[y]=\kappa\},\qquad\operatorname{\mathbf{Sel}}(Y\mid m):=\{y\in\operatorname{\mathbf{Sel}}(Y):Med(y)=m\},$
where $\operatorname*{Med}(y)$ denotes median of $y$. These sets of selections encode additional information about the law of the underlying latent variable $y^{*}$ constrained to lie in $Y(\omega)$ almost surely. Our aim is to understand (i) when such restricted selection sets are nonempty, (ii) show how restricting they can be in terms of the random interval $Y=[y_{L},y_{U}]$ and the underlying probability space, and (iii) which values of basic scalar functionals,
$\mathbb{E}[y],\qquad\operatorname*{Med}(y),\qquad\mathbf{P}(y\in A),$
can be realized by selections $y$ satisfying a given restriction.
The non-atomicity of $(\Omega,\mathcal{A},\mathbf{P})$ allows us to use measurable selection and rearrangement arguments in a sharp way. At a technical level, several of our proofs combine classical properties of the Aumann expectation with Hardy–Littlewood type inequalities and quantile representations. This yields explicit and, in many cases, closed-form descriptions of the admissible ranges of expectations, medians, event probabilities, and higher moments under simple one-dimensional constraints on the selections.
1.1. Related Literature
Our work is rooted in the classical theory of random closed sets and selection expectation. The notion of the Aumann integral for set-valued mappings was introduced in [Aumann1965] and developed further in [Artstein1974], where convexity and support-function representations play a central role. In the one-dimensional setting considered here, the random sets are intervals $Y=[y_{L},y_{U}]$, and the identity $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is a special case of these general results. For the basic theory of random closed sets, measurable selections, and the measurable graph theorem, we refer to [KuratowskiRyll1965, CastaingValadier1977, Molchanov2005].
Rearrangement and quantile methods appear throughout analysis and probability in the study of extremal inequalities. Our arguments for quantile attainability and median-restricted expectations are in the spirit of classical Hardy-Littlewood-Pólya inequalities, as presented for instance in [LiebLoss2001, HardyLittlewoodPolya1952]. We employ these tools in a simple random interval setting, where they yield explicit formulas for the effect of median constraints on the expectation and for extremal event probabilities under a mean restriction.
Random sets and Aumann expectations have been used in a variety of applied contexts, including mathematical economics, finance, and statistics; see, for example, the monograph [Molchanov2005] and the references therein. In particular, models with set-valued or interval-valued observations, or with latent variables constrained to lie in random sets, naturally lead to parameter ranges described in terms of measurable selections. While such applications motivate our interest in restricted selection sets, the present note is purely structural: we work throughout in a one-dimensional probability-theoretic setting and focus on existence, sharpness, and representation results for the restricted ranges of basic scalar functionals.
1.2. Our Contribution
This note develops a detailed description of restricted selection sets for random intervals on a non-atomic probability space. Starting from the benchmark Aumann expectation range and median interval for arbitrary selections $y\in\operatorname{\mathbf{Sel}}(Y)$, we first show that these bounds are sharp in a strong sense. For expectations, every point of $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is realized as $\mathbb{E}[y]$ for some measurable selection (Proposition 2.2). More generally, for each $\alpha\in(0,1)$, any value between the capacity and containment quantiles of $Y$ can be attained as the $\alpha$–quantile of a selection (Proposition 2.4).
We then introduce mean- and median-restricted selection sets and analyze the induced ranges of expectations. In the median-restricted case, we obtain sharp identification intervals for the mean: under mild feasibility conditions, the set
$\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y\mid m)\}$
is a compact interval $\mathbb{E}_{\min}(m),\mathbb{E}_{\max}(m)]$ whose endpoints admit an explicit representation in terms of quantile integrals of the “gaps” $y_{U}-m$ and $m-y_{L}$ on the contact set $\{y_{L}\leq m\leq y_{U}\}$ (Proposition 2.7). The proof relies on two auxiliary rearrangement lemmas, including a quantitative quantile–area identity.
We also examine the converse question of how mean restrictions affect the set of attainable medians. A simple two-state construction (Example 2.11) shows that a mean restriction $\mathbb{E}(y^{*})=\kappa$ does not, in general, reduce the benchmark median interval $\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$: every median in this interval can remain attainable under the mean constraint. For event probabilities $\mathbf{P}(y^{*}\in A)$, we refine the classical random-set bounds in Proposition 2.12 by imposing a mean restriction and deriving a dual (Lagrangian) representation (Theorem 2.13). This dual problem leads to an explicit threshold selection rule (Theorem 2.14) and a closed-form identification formula for $\mathbf{P}(y^{*}\in A)$ in terms of extremal distances between $Y$ and the target set $A$ (Corollary 2.15), while preserving consistency with the unrestricted capacity/containment bounds (Proposition 2.16).
Finally, we indicate extensions of the restricted selection framework in two directions. First, we consider higher-moment constraints and show how to obtain dual representations for the corresponding ranges of expectations using a simple two-dimensional random-set construction (Theorem 2.20). Second, we discuss quantile restrictions at levels $\alpha\neq\tfrac{1}{2}$ and describe the induced intervals of admissible means in this more general setting (Proposition 2.23). These extensions illustrate that the one-dimensional random-interval framework, combined with standard tools from the theory of random sets and rearrangements, already captures a rich collection of sharp identification and extremal results for restricted selections.
1.3. Notation and preliminaries
We work on a fixed, complete, non-atomic probability space
$(\Omega,\mathcal{A},\mathbf{P}).$
All random objects are defined on this space. Expectations are denoted by $\mathbb{E}[\cdot]$, and indicator functions of events $B\in\mathcal{A}$ by $\mathbf{1}_{B}$. For $x\in\mathbb{R}$ we write $x_{+}:=\max\{x,0\}$.
Spaces and sets. We denote by $\mathbb{R}$ the real line and, for $d\in\mathbb{N}$, by $\mathbb{R}^{d}$ the dd–dimensional Euclidean space equipped with its usual topology and Euclidean norm $\|\cdot\|$. Let $\mathcal{K}(\mathbb{R})$ be the family of all nonempty closed subsets of $\mathbb{R}$. Our random sets will always take values in $\mathcal{K}(\mathbb{R})$. For general background on random closed sets we refer to [Molchanov2005].
Random intervals. A random closed set in $\mathbb{R}$ is a measurable map
$Y:(\Omega,\mathcal{A})\to\big(\mathcal{K}(\mathbb{R}),\mathcal{B}(\mathcal{K}(\mathbb{R}))\big),$
where $\mathcal{B}(\mathcal{K}(\mathbb{R}))$ is the Borel $\sigma$–algebra generated by the Fell (hit–or–miss) topology (see [Molchanov2005]). In this note we consider the special case of random intervals, that is,
$Y(\omega)=[y_{L}(\omega),y_{U}(\omega)]\in\mathcal{K}(\mathbb{R}),\qquad\omega\in\Omega,$
for some real-valued random variables $y_{L},y_{U}:\Omega\to\mathbb{R}$ satisfying
$\mathbf{P}\big(y_{L}\leq y_{U}\big)=1.$
Unless otherwise stated we assume
$\mathbb{E}\big(|y_{L}|+|y_{U}|\big)<\infty.$ (1.3)
Selections and Aumann expectation. A (measurable) selection of a random closed set $Y$ is a random variable $y:\Omega\to\mathbb{R}$ such that
$y(\omega)\in Y(\omega)\qquad\text{for $\mathbf{P}$–almost every }\omega.$
The existence of measurable selections for closed-valued measurable multifunctions is classical, see e.g. [KuratowskiRyll1965, CastaingValadier1977]. We denote the family of all measurable selections of $Y$ by
$\operatorname{\mathbf{Sel}}(Y):=\big\{y:\Omega\to\mathbb{R}\ \text{measurable}:y(\omega)\in Y(\omega)\ \text{$\mathbf{P}$–a.s.}\big\},$
and the subfamily of integrable selections by
$\operatorname{\mathbf{Sel}}^{1}(Y):=\big\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}|y|<\infty\big\}.$
Under (1.3), every selection of a random interval $Y=[y_{L},y_{U}]$ is integrable, so $\operatorname{\mathbf{Sel}}(Y)=\operatorname{\mathbf{Sel}}^{1}(Y)$ in this case.
Following [Aumann1965], the Aumann expectation (selection expectation) of a random closed set $Y$ with at least one integrable selection is defined as the closed set
$\mathbb{E}[Y]:=\overline{\big\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}^{1}(Y)\big\}}\subset\mathbb{R},$
where the closure is taken in the usual topology on $\mathbb{R}$; see also [Artstein1974, Molchanov2005]. For a random interval $Y=[y_{L},y_{U}]$ satisfying (1.3), one has the simple representation
$\mathbb{E}[Y]=\big[\mathbb{E}(y_{L}),\,\mathbb{E}(y_{U})\big],$ (1.4)
which is a special case of the general convexity and support–function properties of Aumann integrals.
Capacity and containment functionals. For a random closed set $Y$ and a Borel set $B\subseteq\mathbb{R}$, the capacity (hitting) and containment functionals are defined by
$T_{Y}(B):=\mathbf{P}\big(Y\cap B\neq\emptyset\big),\qquad C_{Y}(B):=\mathbf{P}\big(Y\subseteq B\big),$
see [Molchanov2005]. In the one–dimensional setting we will mainly use the half–line sets $(-\infty,t]$, $t\in\mathbb{R}$, and write
$T_{Y}((-\infty,t])=\mathbf{P}\big(Y\cap(-\infty,t]\neq\emptyset\big),\qquad C_{Y}((-\infty,t])=\mathbf{P}\big(Y\subseteq(-\infty,t]\big).$
Their (left-continuous) generalized inverses are given by
$T_{Y}^{-1}(\alpha):=\inf\{t\in\mathbb{R}:T_{Y}((-\infty,t])\geq\alpha\},\qquad C_{Y}^{-1}(\alpha):=\inf\{t\in\mathbb{R}:C_{Y}((-\infty,t])\geq\alpha\},$
for $\alpha\in(0,1)$. For random intervals $Y=[y_{L},y_{U}]$ one has
$T_{Y}((-\infty,t])=\mathbf{P}(y_{L}\leq t),\qquad C_{Y}((-\infty,t])=\mathbf{P}(y_{U}\leq t),$
so that $T_{Y}^{-1}$ and $C_{Y}^{-1}$ coincide with the quantile functions of $y_{L}$ and $y_{U}$, respectively.
Distribution functions, quantiles, and medians. If $Z$ is a real-valued random variable with distribution function
$F_{Z}(t):=\mathbf{P}(Z\leq t),\qquad t\in\mathbb{R},$
its (left-continuous) quantile function is
$F_{Z}^{-1}(\alpha):=\inf\{t\in\mathbb{R}:F_{Z}(t)\geq\alpha\},\qquad\alpha\in(0,1).$
A (possibly non-unique) median of $Z$ is any $m\in\mathbb{R}$ such that
$\mathbf{P}(Z\leq m)\ \geq\ \tfrac{1}{2},\qquad\mathbf{P}(Z\geq m)\ \geq\ \tfrac{1}{2}.$
We write $\operatorname*{Med}(Z)$ for the (possibly set-valued) set of all such medians; in the scalar case $\operatorname*{Med}(Z)$ is always a nonempty closed interval (possibly degenerate).
Restricted selection sets. The objects of interest in this note are selection sets restricted by simple scalar constraints. For $\kappa\in\mathbb{R}$ we define the $\kappa$-mean restricted selection set by
$\operatorname{\mathbf{Sel}}(Y\mid\kappa):=\big\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}[y]=\kappa\big\},$
and for $m\in\mathbb{R}$ the $m$-median restricted selection set by
$\operatorname{\mathbf{Sel}}(Y\mid m):=\big\{y\in\operatorname{\mathbf{Sel}}(Y):m\in\operatorname*{Med}(y)\big\}.$
2. Main Result
In this section, we study restricted selection sets of a random interval $Y=[y_{L},y_{U}]$ on the non-atomic probability space $(\Omega,\mathcal{A},\mathbf{P})$ introduced in Subsection 1.3. We first record benchmark ranges for the expectation, median, and event probabilities of measurable selections of $Y$, and then investigate how these ranges are sharpened by simple scalar restrictions on the mean or median of the selection. Throughout we assume $y_{L},y_{U}$ are measurable real-valued random variables with $\mathbf{P}(y_{L}\leq y_{U})=1$ and
$\mathbb{E}(|y_{L}|+|y_{U}|)<\infty,$
so that every selection $y\in\operatorname{\mathbf{Sel}}(Y)$ is integrable and $\operatorname{\mathbf{Sel}}(Y)=\operatorname{\mathbf{Sel}}^{1}(Y)$.
2.1. Benchmark identification ranges
Consider a latent scalar random variable $y^{*}$ such that
$\mathbf{P}\big(y_{L}\leq y^{*}\leq y_{U}\big)=1.$
Equivalently, $y^{*}$ is a measurable selection of the random interval $Y=[y_{L},y_{U}]$. Without additional information on the distribution of $y^{*}$, the following benchmark identification ranges for its expectation and median are immediate.
First, by the Aumann expectation identity (1.4),
$\mathbb{E}[Y]=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big],$
so the set of possible expectations of selections equals
$\big\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y)\big\}=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big].$
Thus
$\mathbb{E}(y^{*})\in\Theta_{E}^{*}:=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big].$ (2.1)
Second, since $y_{L}\leq y^{*}\leq y_{U}$ almost surely, the median of $y^{*}$ must lie between the median of $y_{L}$ and the median of $y_{U}$. Writing $\operatorname*{Med}(Z)$ for the (possibly set-valued) set of medians of a real-valued random variable $Z$ (see Subsection 1.3), we obtain
$\operatorname*{Med}(y^{*})\in\Theta_{M}^{*}:=\big[\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})\big],$ (2.2)
where the right-hand side is understood as the smallest interval containing $\operatorname*{Med}(y_{L})$ and $\operatorname*{Med}(y_{U})$.
The bounds in (2.1) and (2.2) are sharp in the sense that every point in these intervals is the expectation (respectively, a median) of some measurable selection $y$ of $Y$. Sharpness for expectations is a direct consequence of (1.4); sharpness for medians will follow from the quantile attainability result in subsection 2.3.
2.2. Restricted selection sets
We now introduce restricted selection sets defined by simple scalar constraints, which will be the central objects of interest.
Definition 2.1 (Mean-restricted selection set).
For $\kappa\in\mathbb{R}$, the $\kappa$-mean restricted selection set is
$\operatorname{\mathbf{Sel}}(Y\mid\kappa):=\big\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}[y]=\kappa\big\},$
and the set $\operatorname{\mathbf{Sel}}^{1}(Y\mid\kappa)$ is defined analogously, replacing $\operatorname{\mathbf{Sel}}(Y)$ by $\operatorname{\mathbf{Sel}}^{1}(Y)$. Under the integrability condition $\mathbb{E}(|y_{L}|+|y_{U}|)<\infty$, these coincide.
Proposition 2.2.
Let $Y=[y_{L},y_{U}]$ be a random interval with measurable endpoints satisfying $y_{L}\leq y_{U}$ almost surely and $y_{L},y_{U}\in L^{1}(\mathbf{P})$. Then for any $\kappa\in[\mathbf{E}(y_{L}),\mathbf{E}(y_{U})]$,
$\operatorname{\mathbf{Sel}}(Y\mid\kappa)\neq\emptyset.$
Proof.
Let $Y=[y_{L},y_{U}]$ be a random interval with measurable endpoints satisfying $y_{L}\leq y_{U}$ a.s. and assume $y_{L},y_{U}\in L^{1}$, the set of finitely integrable random variables. For each $t\in[0,1]$ define
$y_{t}\;\coloneqq\;(1-t)\,y_{L}+t\,y_{U}.$
Then $y_{t}$ is measurable and, since $y_{L}\leq y_{t}\leq y_{U}$ a.s., we have $y_{t}\in\operatorname{\mathbf{Sel}}(Y)$ (indeed $y_{t}\in\operatorname{\mathbf{Sel}}^{1}(Y)$ by integrability). By linearity of expectation,
$\mathbf{E}[y_{t}]\;=\;(1-t)\,\mathbf{E}[y_{L}]+t\,\mathbf{E}[y_{U}],$
so the map $t\mapsto\mathbf{E}[y_{t}]$ is affine with range $\mathbf{E}(y_{L}),\mathbf{E}(y_{U})]$. Given any $\kappa\in[\mathbf{E}(y_{L}),\mathbf{E}(y_{U})]$, choose
$t^{\ast}\;=\;\begin{cases}\dfrac{\kappa-\mathbf{E}(y_{L})}{\mathbf{E}(y_{U})-\mathbf{E}(y_{L})},&\text{if }\mathbf{E}(y_{U})>\mathbf{E}(y_{L}),\\[6.0pt] 0,&\text{if }\mathbf{E}(y_{U})=\mathbf{E}(y_{L}),\end{cases}$
and set $y\coloneqq y_{t^{\ast}}$. Then $y\in\operatorname{\mathbf{Sel}}(Y)$ and $\mathbf{E}[y]=\kappa$, hence $\operatorname{\mathbf{Sel}}(Y\mid\kappa)\neq\emptyset$ (and, under $L^{1}$ endpoints, $\operatorname{\mathbf{Sel}}^{1}(Y\mid\kappa)\neq\emptyset$). ∎
Definition 2.3 (Median-restricted selection set).
For $m\in\mathbb{R}$, the $m$-median restricted selection set is
$\operatorname{\mathbf{Sel}}(Y\mid m):=\big\{y\in\operatorname{\mathbf{Sel}}(Y):m\in\operatorname*{Med}(y)\big\},$
and $\operatorname{\mathbf{Sel}}^{1}(Y\mid m)$ is defined analogously.
Whenever $m$ lies between the capacity and containment medians of $Y$, the non-emptiness of $\operatorname{\mathbf{Sel}}(Y\mid m)$ follows from the quantile attainability result in Subsection 2.3 with $\alpha=\tfrac{1}{2}$.
The rest of this section describes the induced ranges of expectations, medians, and event probabilities under these restrictions.
2.3. Quantile attainability via selections
We now show that any value between the capacity and containment quantiles of $Y$ can be realized as a quantile of some measurable selection. Recall from Subsection 1.3 that for a random closed set $Y$ and $t\in\mathbb{R}$,
$T_{Y}((-\infty,t]):=\mathbf{P}\!\big(Y\cap(-\infty,t]\neq\emptyset\big),\qquad C_{Y}((-\infty,t]):=\mathbf{P}\!\big(Y\subset(-\infty,t]\big),$
and their generalized inverses
$T_{Y}^{-1}(\alpha):=\inf\{t:T_{Y}((-\infty,t])\geq\alpha\},\quad C_{Y}^{-1}(\alpha):=\inf\{t:C_{Y}((-\infty,t])\geq\alpha\},\quad\alpha\in(0,1).$
Proposition 2.4 (Quantile attainability via selections).
Let $(\Omega,\mathcal{F},\mathbf{P})$ be a non-atomic probability space and let $Y$ be a (measurable) random closed convex set in $\mathbb{R}$ (random interval). For $t\in\mathbb{R}$, define
$T_{Y}((-\infty,t]):=\mathbf{P}\!\big(Y\cap(-\infty,t]\neq\emptyset\big),\qquad C_{Y}((-\infty,t]):=\mathbf{P}\!\big(Y\subset(-\infty,t]\big),$
and let the generalized inverses be
$T_{Y}^{-1}(\alpha):=\inf\{t:T_{Y}((-\infty,t])\geq\alpha\},\quad C_{Y}^{-1}(\alpha):=\inf\{t:C_{Y}((-\infty,t])\geq\alpha\},\quad\alpha\in(0,1).$
If $m\in[T_{Y}^{-1}(\alpha),\,C_{Y}^{-1}(\alpha)]$, then there exists a measurable selection $y$ of $Y$ such that $F_{y}^{-1}(\alpha)=m$.
Proof.
Fix $\alpha\in(0,1)$ and $m\in[T_{Y}^{-1}(\alpha),\,C_{Y}^{-1}(\alpha)]$, and write $H_{t}:=(-\infty,t]$. Recall that for an interval-valued random set $Y=[y_{L},y_{U}]$ we have
$T_{Y}(H_{t})=\mathbf{P}(y_{L}\leq t),\qquad C_{Y}(H_{t})=\mathbf{P}(y_{U}\leq t),\qquad t\in\mathbb{R}.$
Partition at $m$. Partition $\Omega$ at the threshold $m$ as
$A^{-}:=\{Y\subset H_{m}\}=\{y_{U}\leq m\},\qquad A^{+}:=\{Y\cap H_{m}=\emptyset\}=\{y_{L}>m\},$
and
$A^{0}:=\Omega\setminus(A^{-}\cup A^{+}).$
Then
$\mathbf{P}(A^{-})=C_{Y}(H_{m}),\qquad\mathbf{P}(A^{-}\cup A^{0})=T_{Y}(H_{m}).$
Since $m\geq T_{Y}^{-1}(\alpha)$, $T_{Y}(H_{m})\geq\alpha$, so
$\mathbf{P}(A^{0})=T_{Y}(H_{m})-\mathbf{P}(A^{-})\ \geq\ \alpha-\mathbf{P}(A^{-}).$
Set
$\delta:=\bigl(\alpha-\mathbf{P}(A^{-})\bigr)_{+}\in[0,\mathbf{P}(A^{0})].$
By non-atomicity there exists a measurable $B\subset A^{0}$ with $\mathbf{P}(B)=\delta$.
Extremal measurable selections. For each $\omega$, define
$Y^{\leq m}(\omega):=Y(\omega)\cap(-\infty,m],\qquad Y^{>m}(\omega):=Y(\omega)\cap(m,\infty).$
On $A^{-}\cup A^{0}$, the set $Y^{\leq m}(\omega)$ is nonempty; on $A^{0}\cup A^{+}$, the set $Y^{>m}(\omega)$ is nonempty. Because $Y$ has a measurable graph and the half-lines are Borel, $\operatorname{graph}(Y^{\leq m})$ and $\operatorname{graph}(Y^{>m})$ are measurable. By standard measurable selection results (e.g. via a Castaing representation, see [KuratowskiRyll1965, CastaingValadier1977]) we can take measurable selections
[…]
with the additional requirement that $z^{-}(\omega)=\sup Y^{\leq m}(\omega)$ and $z^{+}(\omega)=\inf Y^{>m}(\omega)$ on their domains.
Since $Y(\omega)=[y_{L}(\omega),y_{U}(\omega)]$ is an interval, these extremal selections have a simple form:
$z^{-}(\omega)=\begin{cases}y_{U}(\omega),&\omega\in A^{-},\\[2.0pt] m,&\omega\in A^{0},\end{cases}\qquad z^{+}(\omega)=\begin{cases}m,&\omega\in A^{0},\\[2.0pt] y_{L}(\omega),&\omega\in A^{+}.\end{cases}$
In particular,
[…]
For notational convenience we complete these to global selections by setting
$y^{-}(\omega):=\begin{cases}z^{-}(\omega),&\omega\in A^{-}\cup A^{0},\\[2.0pt] y_{L}(\omega),&\omega\in A^{+},\end{cases}\qquad y^{+}(\omega):=\begin{cases}z^{+}(\omega),&\omega\in A^{0}\cup A^{+},\\[2.0pt] y_{U}(\omega),&\omega\in A^{-}.\end{cases}$
Thus $y^{-},y^{+}\in\operatorname{\mathbf{Sel}}(Y)$ and
[…]
Pasting. Define $y:\Omega\to\mathbb{R}$ by
$y(\omega)=\begin{cases}y^{-}(\omega),&\omega\in A^{-}\cup B,\\[2.0pt] y^{+}(\omega),&\omega\in(A^{0}\setminus B)\cup A^{+}.\end{cases}$
Then $y(\omega)\in Y(\omega)$ for all $\omega$, so $y$ is a measurable selection.
On $A^{-}\cup B$ we have $y\leq m$, and on $(A^{0}\setminus B)\cup A^{+}$ we have $y\geq m$, with $y>m$ on $A^{+}$. Hence
$\mathbf{P}(y\leq m)\;\geq\;\mathbf{P}(A^{-})+\mathbf{P}(B)\;=\;\mathbf{P}(A^{-})+\delta\;=\;\max\{\mathbf{P}(A^{-}),\,\alpha\}\;\geq\;\alpha.$
Thus $F_{y}(m):=\mathbf{P}(y\leq m)\geq\alpha$.
Behaviour below $m$. Let $t<m$. On $A^{+}$ and on $A^{0}\setminus B$ we have $y\geq m>t$, so $\{y\leq t\}$ does not occur there. On $B\subset A^{0}$ we have $y(\omega)=y^{-}(\omega)=z^{-}(\omega)=m>t$, so $\{y\leq t\}$ does not occur on $B$ either. On $A^{-}$ we have $y(\omega)=y^{-}(\omega)=z^{-}(\omega)=y_{U}(\omega)$, since $Y(\omega)\subset H_{m}$ implies $Y^{\leq m}(\omega)=Y(\omega)$. Therefore, for $\omega\in A^{-}$,
$y(\omega)\leq t\ \Longrightarrow\ y_{U}(\omega)\leq t\ \Longleftrightarrow\ Y(\omega)\subset H_{t}.$
Consequently,
$\{y\leq t\}\subset\{Y\subset H_{t}\},\qquad\mathbf{P}(y\leq t)\leq\mathbf{P}(Y\subset H_{t})=C_{Y}(H_{t}).$
Because $m\leq C_{Y}^{-1}(\alpha)$ and $t<m$, the monotonicity of $C_{Y}$ implies $C_{Y}(H_{t})<\alpha$, hence
$\mathbf{P}(y\leq t)\ \leq\ C_{Y}(H_{t})\ <\ \alpha,\qquad t<m.$
Identify $F_{y}^{-1}(\alpha)$. We have shown that $\mathbf{P}(y\leq t)<\alpha$ for all $t<m$, while $\mathbf{P}(y\leq m)\geq\alpha$. By the definition of the generalized inverse,
$F_{y}^{-1}(\alpha)=\inf\{t:\mathbf{P}(y\leq t)\geq\alpha\}=m.$
This completes the proof. ∎
Taking $\alpha=\tfrac{1}{2}$, the previous proposition implies that every point in $\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$ is a median of some selection of $Y$, so the benchmark median interval $\Theta_{M}^{*}$ is sharp.
2.4. Auxiliary rearrangement lemmas
We next record two auxiliary results that will be used to characterize the effect of median restrictions on expectations.
Lemma 2.5 (Conditional rearrangement / quantile bound[LiebLoss2001]).
Let $(\Omega,\mathcal{F},\mathbf{P})$ be a probability space, let $M\in\mathcal{F}$ with $p_{0}:=\mathbf{P}(M)>0$, and let $X:\Omega\to[0,\infty)$ be $\mathcal{F}$-measurable. Denote by $\mathcal{F}_{M}:=\{A\cap M:\ A\in\mathcal{F}\}$ the trace $\sigma$-algebra on $M$, and by
$\mathbf{P}_{M}(\,\cdot\,):=\mathbf{P}(\,\cdot\,\mid M)=\frac{\mathbf{P}((\,\cdot\,)\cap M)}{p_{0}}$
the conditional probability on $(M,\mathcal{F}_{M})$. Write $F_{X\mid M}(t):=\mathbf{P}_{M}(X\leq t)$ for the conditional distribution of $X$ given $M$ and let
$Q(u):=F^{-1}_{X\mid M}(u):=\inf\{t\in\mathbb{R}:\ F_{X\mid M}(t)\geq u\},\qquad u\in[0,1],$
be the (left-continuous) conditional quantile.
Then, (1) for any $S\in\mathcal{F}_{M}$ with $\mathbf{P}(S)=s\in[0,p_{0}]$,
$\mathbb{E}\big[X\,\mathbf{1}_{S}\big]\ \geq\ p_{0}\int_{0}^{s/p_{0}}Q(u)\,du.$ (2.3)
(2) If the probability space $(M,\mathcal{F}_{M},\mathbf{P}_{M})$ is atomless, equality in (2.3) is attained by any least-$X$ set $S$ of mass $s$, i.e.
$S=\{X<q\}\cap M\;\cup\;\Big(\{X=q\}\cap M\cap B\Big),$
where $q:=Q(s/p_{0})$ and $B\in\mathcal{F}_{M}$ is chosen with $\mathbf{P}_{M}(B)=\theta$ to adjust the measure when $\mathbf{P}_{M}(X=q)>0$.
(3) Without the atomlessness assumption, equality is attained on an atomless product extension by splitting atoms with an auxiliary independent uniform variable.
Proof.
Step 0 (Reduction to the conditional space). Since $\mathbf{P}(S)=s$ if and only if $\mathbf{P}_{M}(S)=\alpha:=s/p_{0}$, and
$\mathbb{E}\big[X\,\mathbf{1}_{S}\big]\;=\;p_{0}\,\mathbb{E}_{M}\big[X\,\mathbf{1}_{S}\big],$
it suffices to prove that, on the probability space $(M,\mathcal{F}_{M},\mathbf{P}_{M})$,
$\inf_{\begin{subarray}{c}S\in\mathcal{F}_{M}\\ \mathbf{P}_{M}(S)=\alpha\end{subarray}}\mathbb{E}_{M}[X\,\mathbf{1}_{S}]\;=\;\int_{0}^{\alpha}Q(u)\,du,\qquad\alpha\in[0,1].$ (2.4)
Step 1 (Distributional transform and quantile representation). If $(M,\mathcal{F}_{M},\mathbf{P}_{M})$ is not atomless, pass to the product extension
$(\bar{\Omega},\bar{\mathcal{F}},\bar{\mathbf{P}})\;:=\;\big(M\times[0,1],\ \mathcal{F}_{M}\otimes\mathcal{B}([0,1]),\ \mathbf{P}_{M}\otimes\lambda\big),$
where $\lambda$ is Lebesgue measure and let $V:\bar{\Omega}\to[0,1]$ be the canonical uniform variable. Define on $(\bar{\Omega},\bar{\mathcal{F}},\bar{\mathbf{P}})$ the distributional transform
$U\ :=\ F_{X\mid M}(X-)\ +\ V\cdot\Delta F_{X\mid M}(X),\qquad\text{where}\quad\Delta F_{X\mid M}(x):=F_{X\mid M}(x)-\lim_{t\uparrow x}F_{X\mid M}(t).$
Then $U\sim\mathrm{Unif}[0,1]$ (under $\mathbf{P}_{M}$ or $\bar{\mathbf{P}}$, as appropriate) and
$X\ =\ Q(U)\qquad\text{$\mathbf{P}_{M}$-a.s.\ (or $\bar{\mathbf{P}}$-a.s.)}.$ (2.5)
(When $F_{X\mid M}$ is continuous, one may simply take $U=F_{X\mid M}(X)$ on $M$.)
Step 2 (From sets to weights depending only on $U$). Fix any $S\in\mathcal{F}_{M}$ with $\mathbf{P}_{M}(S)=\alpha$ and set $H:=\mathbf{1}_{S}$. Using (2.5) and the tower property,
$\mathbb{E}_{M}[X\,H]\;=\;\mathbb{E}_{M}\!\big[Q(U)\,H\big]\;=\;\mathbb{E}_{M}\!\Big[\mathbb{E}_{M}\!\big[Q(U)\,H\,\big|\,U\big]\Big]\;=\;\mathbb{E}_{M}\!\big[Q(U)\,h(U)\big],$
where
[…]
because $U\sim\mathrm{Unif}[0,1]$. Thus the minimization in (2.4) reduces to the deterministic calculus of variations problem
$\inf\Big\{\ \int_{0}^{1}Q(u)\{h(u)\,du\ :\ 0\leq h\leq 1,\ \int_{0}^{1}h=\alpha\ \Big\}.$ (2.6)
Here $Q$ is nondecreasing and integrable on $[0,1]$ (since $X\geq 0$ is integrable under $\mathbf{P}_{M}$).
Step 3 (Bathtub principle: optimal $h$ is a threshold). We show that any minimizer of (2.6) is of the form
$h^{*}(u)=\mathbf{1}_{\{Q(u)<\lambda\}}\;+\;\theta\,\mathbf{1}_{\{Q(u)=\lambda\}},$
for some $\lambda\in\mathbb{R}$ and $\theta\in[0,1]$ chosen to satisfy $\int_{0}^{1}h^{*}=\alpha$. The argument is standard:
Suppose $h$ is feasible and not of this form. Then there exist $u_{1},u_{2}$ with $Q(u_{1})<Q(u_{2})$, $h(u_{1})<1$, and $h(u_{2})>0$. For $\varepsilon>0$ small, define
$\tilde{h}:=h+\varepsilon\,\mathbf{1}_{B_{1}}-\varepsilon\,\mathbf{1}_{B_{2}},$
where $B_{1}\subset\{u:\ Q(u)\leq Q(u_{1})\}$ and $B_{2}\subset\{u:\ Q(u)\geq Q(u_{2})\}$ are measurable subsets of equal Lebesgue measure chosen so that $0\leq\tilde{h}\leq 1$ a.e. and $\int\tilde{h}=\int h=\alpha$ (possible because $h(u_{1})<1$ and $h(u_{2})>0$ on sets of positive measure). Then
$\int_{0}^{1}Q(u)\,\tilde{h}(u)\,du\;=\;\int_{0}^{1}Q(u)\,h(u)\,du\;+\;\varepsilon\!\left(\int_{B_{1}}Q(u)\,du-\int_{B_{2}}Q(u)\,du\right)$
$\int_{0}^{1}Q(u)\,h(u)\,du,$
since $Q$ is nondecreasing and $Q(u_{1})<Q(u_{2})$. Hence $h$ is not optimal. Iterating this exchange shows that any optimal $h$ must take values $1$ on the region where $Q$ is smallest, 0 where $Q$ is largest, and possibly a fractional value on a single level set $\{Q=\lambda\}$ to match the total mass $\alpha$. Because $Q$ is nondecreasing, such an $h^{*}$ is equivalent (modulo null sets) to the simple threshold
$h^{*}(u)=\mathbf{1}_{[0,\alpha]}(u),$
[…]
Therefore,
[…]
Step 4 (Attainment by a set $S$). If $(M,\mathcal{F}_{M},\mathbf{P}_{M})$ is atomless, there exists $B\in\mathcal{F}_{M}$ with any prescribed $\mathbf{P}_{M}(B)\in[0,1]$. Let $q:=Q(\alpha)$ and choose $B\subseteq\{X=q\}\cap M$ with $\mathbf{P}_{M}(B)=\theta$, where $\theta$ is chosen so that
$\mathbf{P}_{M}\big(\{X<q\}\cap M\big)+\theta\,\mathbf{P}_{M}(\{X=q\}\cap M)=\alpha.$
Define
$S^{*}\;:=\;\big(\{X<q\}\cap M\big)\ \cup\ B.$
Then $\mathbf{P}_{M}(S^{*})=\alpha$, and $H^{*}:=\mathbf{1}_{S^{*}}$ satisfies $H^{*}=h^{*}(U)$ in the sense of Step 3; hence
$\mathbb{E}_{M}[X\,\mathbf{1}_{S^{*}}]\;=\;\int_{0}^{\alpha}Q(u)\,du,$
establishing equality in (2.4).
If $(M,\mathcal{F}_{M},\mathbf{P}_{M})$ has atoms on $\{X=q\}$ and no suitable $B\in\mathcal{F}_{M}$ exists with the exact conditional mass, pass to the product extension of Step 1. Use the auxiliary uniform $V$ to split the level set measurably: set
$S^{*}\;:=\;\big(\{X<q\}\cap M\big)\ \cup\ \Big(\{X=q\}\cap M\cap\{V\leq\theta\}\Big),$
now a set in $\bar{\mathcal{F}}$ with $\bar{\mathbf{P}}(S^{*})=\alpha$ and $\mathbb{E}_{\bar{\mathbf{P}}}[X\,\mathbf{1}_{S^{*}}]=\int_{0}^{\alpha}Q(u)\,du$ (note $X$ does not depend on $V$). This yields equality on the extension.
Step 5. Combining Steps 1–4, we have proved (2.4). Multiplying by $p_{0}$ gives (2.3). The stated forms of the equality sets follow from the constructions above.
Edge cases. If $\alpha=0$ (i.e., $s=0$), both sides vanish. If $\alpha=1$ (i.e., $s=p_{0}$), both sides equal $\mathbb{E}[X\,\mathbf{1}_{M}]=p_{0}\int_{0}^{1}Q(u)\,du$. ∎
Lemma 2.6 (Quantile–area identity [HardyLittlewoodPolya1952]).
Let $Z$ be a nonnegative real-valued random variable with distribution function $F_{Z}$ and (left-continuous) quantile function
$F_{Z}^{-1}(u)\;:=\;\inf\{t\in\mathbb{R}:F_{Z}(t)\geq u\},\qquad u\in(0,1).$
Then for every $\alpha\in(0,1)$,
$\int_{0}^{\alpha}F_{Z}^{-1}(u)\,du\;=\;\int_{0}^{\infty}(\alpha-F_{Z}(t))_{+}\,dt,$ (2.8)
where $(x)_{+}:=\max\{x,0\}$.
Proof.
Set $Q(u):=F_{Z}^{-1}(u)$ for $u\in(0,1)$ and fix $\alpha\in(0,1)$. Since $Z\geq 0$ almost surely, we have $Q(u)\geq 0$ for all $u\in(0,1)$. Using the layer-cake representation and Tonelli’s theorem,
$\int_{0}^{\alpha}Q(u)\,du$ $=\int_{0}^{\alpha}\left(\int_{0}^{Q(u)}dt\right)du$
$=\int_{0}^{\infty}\left(\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du\right)dt.$
Thus, it suffices to compute, for each fixed $t\geq 0$, the inner integral
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du$
as a function of $t$.
By definition of the quantile function,
$Q(u)\leq t\quad\Longleftrightarrow\quad u\leq F_{Z}(t),$
because $F_{Z}(t)=\sup\{u\in[0,1]:Q(u)\leq t\}$ for each $t$. Hence
$t\leq Q(u)\quad\Longleftrightarrow\quad u\geq F_{Z}(t).$
Therefore, for fixed $t\geq 0$,
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du=\lambda\bigl(\{u\in[0,\alpha]:u\geq F_{Z}(t)\}\bigr),$
where $\lambda$ denotes Lebesgue measure on $\mathbb{R}$. This set is empty if $F_{Z}(t)\geq\alpha$, and it is the interval $F_{Z}(t),\alpha]$ if $F_{Z}(t)<\alpha$. Consequently,
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du=\begin{cases}\alpha-F_{Z}(t),&\text{if }F_{Z}(t)<\alpha,\\[3.0pt] 0,&\text{if }F_{Z}(t)\geq\alpha,\end{cases}$
which is exactly $(\alpha-F_{Z}(t))_{+}$. Plugging this back into the layer-cake formula gives
$\int_{0}^{\alpha}Q(u)\,du=\int_{0}^{\infty}(\alpha-F_{Z}(t))_{+}\,dt,$
which is (2.8). ∎
2.5. Identification of $\mathbb{E}(y^{*})$ under a median restriction
We now consider the effect of knowing a median on the range of admissible expectations. Let $Y=[y_{L},y_{U}]$ be a random interval with $\mathbb{E}|y_{L}|+\mathbb{E}|y_{U}|<\infty$, and suppose we know that some selection $y\in\operatorname{\mathbf{Sel}}(Y)$ has population median $m\in\mathbb{R}$ in the sense that
$\mathbf{P}(y\leq m)\geq\tfrac{1}{2},\qquad\mathbf{P}(y\geq m)\geq\tfrac{1}{2}.$
We are interested in the range
$\big\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y\mid m)\big\}.$
Define
$A_{-}:=\{y_{U}<m\},\qquad A_{+}:=\{y_{L}>m\},\qquad M:=\{y_{L}\leq m\leq y_{U}\},$
with $p_{-}:=\mathbf{P}(A_{-})$, $p_{+}:=\mathbf{P}(A_{+})$, $p_{0}:=\mathbf{P}(M)=1-p_{-}-p_{+}$. Set the required “mass shortfalls”
$\alpha_{-}(m):=\big(\tfrac{1}{2}-p_{-}\big)_{+},\qquad\alpha_{+}(m):=\big(\tfrac{1}{2}-p_{+}\big)_{+},$
and on $M$ define the nonnegative gaps to the median
$U_{m}:=y_{U}-m,\qquad L_{m}:=m-y_{L}.$
Let $F^{-1}_{U_{m}\mid M}$ and $F^{-1}_{L_{m}\mid M}$ denote the (left-continuous) quantile functions of the conditional laws of $U_{m}$ and $L_{m}$ given $M$.
Proposition 2.7 (Median-restricted Aumann expectation).
Assume feasibility, $p_{-}\leq\tfrac{1}{2}$ and $p_{+}\leq\tfrac{1}{2}$. Then the identified set of means under the median restriction
$\mathcal{S}_{m}:=\{\,y\in\operatorname{\mathbf{Sel}}(Y):\ \operatorname*{Med}(y)=m\,\}$
is the closed interval
$\mathbb{E}[Y\mid\mathcal{S}_{m}]\;=\;\Big[\,\mathbb{E}_{\min}(m),\ \mathbb{E}_{\max}(m)\,\Big],$ (2.9)
with endpoints
$\mathbb{E}_{\max}(m)\;=\;\mathbb{E}[y_{U}]\;-\;p_{0}\int_{0}^{\alpha_{-}(m)/p_{0}}\!F^{-1}_{\,U_{m}\mid M}(u)\,du,$ (2.10)
and
$\mathbb{E}_{\min}(m)\;=\;\mathbb{E}[y_{L}]\;+\;p_{0}\int_{0}^{\alpha_{+}(m)/p_{0}}\!F^{-1}_{\,L_{m}\mid M}(u)\,du,$ (2.11)
where the integrals are understood as 0 when $p_{0}=0$. Both endpoints are attained by measurable selections, and every value in $\mathbb{E}_{\min}(m),\mathbb{E}_{\max}(m)]$ is attainable.
Proof.
Recall
$A_{-}:=\{y_{U}<m\},\quad A_{+}:=\{y_{L}>m\},\quad M:=\{y_{L}\leq m\leq y_{U}\},$
with $p_{-}:=\mathbf{P}(A_{-})$, $p_{+}:=\mathbf{P}(A_{+})$, $p_{0}:=\mathbf{P}(M)$, the shortfalls
$\alpha_{-}(m):=(\tfrac{1}{2}-p_{-})_{+},\qquad\alpha_{+}(m):=(\tfrac{1}{2}-p_{+})_{+},$
and the nonnegative gaps on $M$,
$U_{m}:=y_{U}-m,\qquad L_{m}:=m-y_{L}.$
We assume feasibility $p_{-}\leq\tfrac{1}{2}$ and $p_{+}\leq\tfrac{1}{2}$, which implies $\alpha_{-}(m)+\alpha_{+}(m)=p_{0}$ and ensures that the median constraint can be satisfied by allocating on $M$ a mass $\alpha_{-}(m)$ at or below $m$ and a mass $\alpha_{+}(m)$ at or above $m$.
Upper endpoint. Let $y\in\operatorname{\mathbf{Sel}}(Y)$ satisfy the median restriction $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$. Since $y\leq m$ on $A_{-}$ and $y\leq y_{U}$ on $M$, we must have
$\mathbf{P}\big(\{y\leq m\}\cap M\big)\ \geq\ \tfrac{1}{2}-\mathbf{P}(A_{-})\ =\ \alpha_{-}(m).$
On $\{y\leq m\}\cap M$, $y_{U}-y\geq U_{m}$. Therefore
$\mathbb{E}[y_{U}-y]\ \geq\ \mathbb{E}\!\big[U_{m}\,\mathbf{1}_{\{y\leq m\}\cap M}\big]\ \geq\ \inf_{S\subset M:\ \mathbf{P}(S)\geq\alpha_{-}(m)}\mathbb{E}\!\big[U_{m}\,\mathbf{1}_{S}\big].$
By Lemma 2.5 applied to $X=U_{m}\geq 0$ on $M$,
$\inf_{S\subset M:\ \mathbf{P}(S)\geq\alpha_{-}(m)}\mathbb{E}\!\big[U_{m}\,\mathbf{1}_{S}\big]\ =\ p_{0}\int_{0}^{\alpha_{-}(m)/p_{0}}F^{-1}_{U_{m}\mid M}(u)\,du.$
Consequently,
$\mathbb{E}[y]\ \leq\ \mathbb{E}[y_{U}]-\mathbb{E}[y_{U}-y]\ \leq\ \mathbb{E}[y_{U}]\;-\;p_{0}\int_{0}^{\alpha_{-}(m)/p_{0}}F^{-1}_{U_{m}\mid M}(u)\,du\ :=\ \mathbb{E}_{\max}(m).$
Lower endpoint. Symmetrically, the median restriction implies
$\mathbf{P}\big(\{y\geq m\}\cap M\big)\ \geq\ \tfrac{1}{2}-\mathbf{P}(A_{+})\ =\ \alpha_{+}(m),$
and on $\{y\geq m\}\cap M$ one has $y-y_{L}\geq L_{m}$. Thus
$\mathbb{E}[y-y_{L}]\ \geq\ \inf_{S\subset M:\ \mathbf{P}(S)\geq\alpha_{+}(m)}\mathbb{E}\!\big[L_{m}\,\mathbf{1}_{S}\big]\ =\ p_{0}\int_{0}^{\alpha_{+}(m)/p_{0}}F^{-1}_{L_{m}\mid M}(u)\,du,$
by Lemma 2.5, which yields
$\mathbb{E}[y]\ \geq\ \mathbb{E}[y_{L}]\;+\;p_{0}\int_{0}^{\alpha_{+}(m)/p_{0}}F^{-1}_{L_{m}\mid M}(u)\,du\ :=\ \mathbb{E}_{\min}(m).$
Attainment. Let $q_{-}:=F^{-1}_{U_{m}\mid M}(\alpha_{-}(m)/p_{0})$ and choose a measurable $S_{-}\subset M$ with $\mathbf{P}(S_{-})=\alpha_{-}(m)$ such that $U_{m}$ is $\mathbf{P}$-a.s. minimal on $S_{-}$, i.e. $S_{-}=\{U_{m}<q_{-}\}\cup(\{U_{m}=q_{-}\}\cap B_{-})$ with suitable measurable $B_{-}$. Define
$\bar{y}(\omega):=\begin{cases}m(\omega),&\omega\in S_{-},\\ y_{U}(\omega),&\omega\in M\setminus S_{-},\\ y_{U}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\bar{y}\in\operatorname{\mathbf{Sel}}(Y)$, $\mathbf{P}(\bar{y}\leq m)=p_{-}+\alpha_{-}(m)=\tfrac{1}{2}$, $\mathbf{P}(\bar{y}\geq m)=\tfrac{1}{2}$, and by construction
$\mathbb{E}[y_{U}-\bar{y}]\ =\ \mathbb{E}\!\big[U_{m}\,\mathbf{1}_{S_{-}}\big]\ =\ p_{0}\int_{0}^{\alpha_{-}(m)/p_{0}}F^{-1}_{U_{m}\mid M}(u)\,du,$
so $\mathbb{E}[\bar{y}]=\mathbb{E}_{\max}(m)$.
Likewise, let $q_{+}:=F^{-1}_{L_{m}\mid M}(\alpha_{+}(m)/p_{0})$ and choose a measurable $S_{+}\subset M$ of measure $\alpha_{+}(m)$ which minimizes $L_{m}$. Define
$\underline{y}(\omega):=\begin{cases}m(\omega),&\omega\in S_{+},\\ y_{L}(\omega),&\omega\in M\setminus S_{+},\\ y_{L}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\underline{y}\in\operatorname{\mathbf{Sel}}(Y)$, $\mathbf{P}(\underline{y}\geq m)=p_{+}+\alpha_{+}(m)=\tfrac{1}{2}$, $\mathbf{P}(\underline{y}\leq m)=\tfrac{1}{2}$, and
$\mathbb{E}[\underline{y}-y_{L}]\ =\ \mathbb{E}\!\big[L_{m}\,\mathbf{1}_{S_{+}}\big]\ =\ p_{0}\int_{0}^{\alpha_{+}(m)/p_{0}}F^{-1}_{L_{m}\mid M}(u)\,du,$
so $\mathbb{E}[\underline{y}]=\mathbb{E}_{\min}(m)$.
Filled interval. For $\theta\in[0,1]$ put $y_{\theta}:=\theta\,\bar{y}+(1-\theta)\,\underline{y}$. Since $Y(\omega)$ is an interval, $y_{\theta}(\omega)\in Y(\omega)$ a.s., hence $y_{\theta}\in\operatorname{\mathbf{Sel}}(Y)$. Moreover,
$\{y_{\theta}\leq m\}\supseteq A_{-}\cup S_{-},\qquad\{y_{\theta}\geq m\}\supseteq A_{+}\cup S_{+},$
so $\mathbf{P}(y_{\theta}\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y_{\theta}\geq m)\geq\tfrac{1}{2}$; thus $y_{\theta}\in\mathcal{S}_{m}$. Linearity of expectation gives
$\mathbb{E}[y_{\theta}]=\theta\,\mathbb{E}_{\max}(m)+(1-\theta)\,\mathbb{E}_{\min}(m),$
so every value in $\mathbb{E}_{\min}(m),\mathbb{E}_{\max}(m)]$ is attained. This establishes (2.9)–(2.10). ∎
Remark 2.8 (Remarks on edge cases).
If $p_{0}=0$ and feasibility holds, one must have $p_{-}=p_{+}=\tfrac{1}{2}$, the integrals in (2.10) vanish, and the median restriction does not tighten $\mathbb{E}[Y]=[\mathbb{E}y_{L},\mathbb{E}y_{U}]$. If $m\in[y_{L},y_{U}]$ a.s., then $p_{-}=p_{+}=0$, $\alpha_{-}(m)=\alpha_{+}(m)=\tfrac{1}{2}$, and both quantile integrals run to $1/2$, recovering the familiar constant-threshold formulas (e.g., for $Y=[a,b]$ a.s., the interval $\tfrac{a+m}{2},\tfrac{m+b}{2}]$).
Example 2.9 (Median-restricted Aumann expectation).
To illustrate Proposition 2.7, consider a stylized selection model
$Y=[y_{L},y_{U}],$
where $y_{L}$ and $y_{U}$ are constructed as follows. Let $U\sim\mathrm{Unif}(0,1)$ and write $F_{L},F_{U}$ for the distribution functions of $\chi^{2}_{(2)}$ and $\chi^{2}_{(5)}$. Set
$y_{L}:=F_{L}^{-1}(U),\qquad y_{U}:=F_{U}^{-1}(U).$
Then $y_{L}\sim\chi^{2}_{(2)}$, $y_{U}\sim\chi^{2}_{(5)}$, and $y_{L}\leq y_{U}$ almost surely.
The medians of the lower and upper bounds are
$m_{L}:=\operatorname*{Med}(y_{L})=1.386,\qquad m_{U}:=\operatorname*{Med}(y_{U})=4.351,$
and we fix a target median
$m:=0.3\,m_{L}+0.7\,m_{U}=3.46.$
We interpret $m$ as the known population median of a measurable selection $y\in\operatorname{\mathbf{Sel}}(Y)$, in the sense that $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y\geq m)\geq\tfrac{1}{2}$.
In this configuration the identification region for the mean simplifies to
$\mathbb{E}(y^{*})\in\big[\,\mathbb{E}(y_{L})+S_{L},\ \mathbb{E}(y_{U})-S_{U}\,\big],$
where
$S_{L}=\int_{m_{L}}^{m}\!\big(F_{L}(t)-\tfrac{1}{2}\big)\,dt,\qquad S_{U}=\int_{m}^{m_{U}}\!\big(\tfrac{1}{2}-F_{U}(t)\big)\,dt.$
The extremal CDFs that achieve the endpoints $\mathbb{E}_{\min}(m)$ and $\mathbb{E}_{\max}(m)$ in (2.9)–(2.10) are obtained by locally reshaping $F_{L}$ on $m_{L},m]$ and $F_{U}$ on $m,m_{U}]$ so that exactly one half of the mass of the induced selection lies below and above $m$, respectively. This construction is visualized in Figure 1.
Remark 2.10 (Derivation of $S_{L}$ and $S_{U}$).
For completeness, we show how the expressions for $S_{L}$ and $S_{U}$ in Example 2.9 arise from the general formulas in Proposition 2.7.
A useful identity is that for any nonnegative real-valued random variable $Z$ with CDF $F_{Z}$ and any $\alpha\in(0,1)$ by Lemma 2.6
$\int_{0}^{\alpha}F_{Z}^{-1}(u)\,du\;=\;\int_{0}^{\infty}(\alpha-F_{Z}(t))_{+}\,dt,$ (2.12)
where $(x)_{+}:=\max\{x,0\}$. This follows from a standard layer-cake (rearrangement) argument.
In the notation of Proposition 2.7, take
$L_{m}:=m-y_{L},\qquad U_{m}:=y_{U}-m,$
so that $L_{m},U_{m}\geq 0$ and, writing $F_{L},F_{U}$ for the CDFs of $y_{L},y_{U}$,
$F_{L_{m}}(\ell)=\mathbf{P}(L_{m}\leq\ell)=\mathbf{P}(y_{L}\geq m-\ell)=1-F_{L}(m-\ell),$
$F_{U_{m}}(\ell)=\mathbf{P}(U_{m}\leq\ell)=\mathbf{P}(y_{U}\leq m+\ell)=F_{U}(m+\ell).$
In the schematic configuration we have $m_{L}<m<m_{U}$, so the median shortfalls are symmetric and Proposition 2.7 reduces to
$\mathbb{E}_{\min}(m)=\mathbb{E}[y_{L}]+\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du,\qquad\mathbb{E}_{\max}(m)=\mathbb{E}[y_{U}]-\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du.$
Apply (2.12) with $Z=L_{m}$ and $\alpha=\tfrac{1}{2}$:
$\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{L_{m}}(\ell)\bigr)_{+}\,d\ell=\int_{0}^{\infty}\bigl(F_{L}(m-\ell)-\tfrac{1}{2}\bigr)_{+}\,d\ell.$
With the change of variable $t=m$-$\ell$ and $d\ell=-dt$, this becomes
$\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du=\int_{-\infty}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)_{+}\,dt=\int_{m_{L}}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)\,dt,$
since $F_{L}(t)\leq\tfrac{1}{2}$ for $t\leq m_{L}$ and the integrand is zero there. This yields
$\mathbb{E}_{\min}(m)=\mathbb{E}[y_{L}]+\int_{m_{L}}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)\,dt=\mathbb{E}[y_{L}]+S_{L}.$
Similarly, with $Z=U_{m}$ and $\alpha=\tfrac{1}{2}$,
$\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{U_{m}}(\ell)\bigr)_{+}\,d\ell=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{U}(m+\ell)\bigr)_{+}\,d\ell.$
Setting $t=m+\ell$ gives
$\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du=\int_{m}^{\infty}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)_{+}\,dt=\int_{m}^{m_{U}}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)\,dt,$
because $F_{U}(t)\geq\tfrac{1}{2}$ for $t\geq m_{U}$. Hence
$\mathbb{E}_{\max}(m)=\mathbb{E}[y_{U}]-\int_{m}^{m_{U}}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)\,dt=\mathbb{E}[y_{U}]-S_{U}.$
This shows how the “cost” terms $S_{L}$ and $S_{U}$ in Example 2.9 arise from the general quantile-integral representation in Proposition 2.7 by a simple change of variables.
Figure 1. Schematic construction of the extremal boundary distributions in Example 2.9. Thin black and gray curves show the baseline CDFs $F_{L}$ and $F_{U}$ of the chi-square lower and upper bounds. The bold green (resp. red) segments indicate the local modification of $F_{L}$ on $m_{L},m]$ (resp. of $F_{U}$ on $m,m_{U}]$) required to satisfy the median constraints $\mathbf{P}(y\leq m)=\mathbf{P}(y\geq m)=\tfrac{1}{2}$ for an underlying selection $y\in\operatorname{\mathbf{Sel}}(Y)$. Vertical lines at $_{L}, mm$, and $m_{U}$ and the bullets at height $1/2$ highlight the binding median restrictions.
2.6. Identification of $\operatorname*{Med}(y^{*})$ under a mean restriction
We now consider the reverse question: given a mean restriction $\mathbb{E}(y^{*})=\kappa$, does this necessarily sharpen the benchmark median interval $\Theta_{M}^{*}=[\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$? The next example shows that the answer is negative in general.
Example 2.11 (Two–state random interval with symmetry around $\kappa$).
Fix $\kappa\in\mathbb{R}$ and $d>0$. On a probability space with two equiprobable states $\{\omega_{1},\omega_{2}\}$, define the random interval
$Y(\omega_{1})=[\,\kappa-2d,\ \kappa\,],\qquad Y(\omega_{2})=[\,\kappa,\ \kappa+2d\,],$
equivalently
$(y_{L},y_{U})=\begin{cases}(\kappa-2d,\ \kappa),&\omega=\omega_{1},\\ (\kappa,\ \kappa+2d),&\omega=\omega_{2},\end{cases}\qquad\mathbf{P}(\omega_{1})=\mathbf{P}(\omega_{2})=\tfrac{1}{2}.$
Then $y_{L}$ takes values $\{\kappa-2d,\kappa\}$ and $y_{U}$ takes values $\{\kappa,\kappa+2d\}$ (each with probability $1/2$), so
$\operatorname*{Med}(y_{L})=\kappa-2d,\qquad\operatorname*{Med}(y_{U})=\kappa,$
and the unrestricted median bound is $\operatorname*{Med}(y^{*})\in[\kappa-2d,\kappa]$.
Now fix any $m\in[\kappa-2d,\kappa]$ and define the selection
$y(\omega_{1}):=m,\qquad y(\omega_{2}):=2\kappa-m.$
Because $m\in[\kappa-2d,\kappa]$ and $2\kappa-m\in[\kappa,\kappa+2d]$, we have $y(\omega_{i})\in Y(\omega_{i})$ for $i=1,2$, hence $y\in\operatorname{\mathbf{Sel}}(Y)$. Moreover,
$\mathbb{E}[y]=\tfrac{1}{2}\,m+\tfrac{1}{2}\,(2\kappa-m)=\kappa,$
so $y$ satisfies the mean restriction. Since the distribution of $y$ places probability $1/2$ at $m$ and $1/2$ at $2\kappa-m$ with $m\leq\kappa\leq 2\kappa-m$, it follows that
$\mathbf{P}(y\leq m)=\tfrac{1}{2},\qquad\mathbf{P}(y\geq m)=1,$
hence $\operatorname*{Med}(y)=m$ (under the standard definition $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y\geq m)\geq\tfrac{1}{2}$).
Because $m$ was arbitrary in $\kappa-2d,\kappa]$, we conclude
$\big\{\operatorname*{Med}(y):\,y\in\operatorname{\mathbf{Sel}}(Y),\ \mathbb{E}[y]=\kappa\big\}=[\kappa-2d,\ \kappa]=\big[\operatorname*{Med}(y_{L}),\,\operatorname*{Med}(y_{U})\big].$
Thus, in this random–interval design, knowing $\mathbb{E}(y^{*})=\kappa$ does not shrink the identified set for $\operatorname*{Med}(y^{*})$.
2.7. Event probabilities: unrestricted and mean-restricted ranges
We now turn to the range of probabilities $\mathbf{P}(y^{*}\in A)$ as $y^{*}$ varies over selections of $Y$, with and without a mean restriction. Let $A$ be a fixed Borel subset of $\mathbb{R}$.
Proposition 2.12 (Unrestricted random-set bounds).
Let $A$ be a closed subset of $\mathbb{R}$. Then for any $y\in\operatorname{\mathbf{Sel}}(Y)$,
$\mathbf{P}\!\big(Y\subseteq A\big)\ \leq\ \mathbf{P}(y\in A)\ \leq\ \mathbf{P}\!\big(Y\cap A\neq\varnothing\big).$ (2.13)
Both bounds are sharp in the sense that there exist measurable selections whose probabilities are arbitrarily close to the respective endpoints.
Proof.
If $Y(\omega)\subseteq A$ and $y(\omega)\in Y(\omega)$, then $y(\omega)\in A$; thus $\{Y\subseteq A\}\subseteq\{y\in A\}$ and $\mathbf{P}(y\in A)\geq\mathbf{P}(Y\subseteq A)$. Similarly, if $y(\omega)\in A$ then $Y(\omega)\cap A\neq\varnothing$, so $\{y\in A\}\subseteq\{Y\cap A\neq\varnothing\}$ and $\mathbf{P}(y\in A)\leq\mathbf{P}(Y\cap A\neq\varnothing)$.
For sharpness, define
$A_{1}:=\{\omega:Y(\omega)\subseteq A\},\qquad A_{2}:=\{\omega:Y(\omega)\cap A\neq\varnothing\}.$
For the upper bound, define a set-valued map
$G_{+}(\omega):=\begin{cases}Y(\omega)\cap A,&\omega\in A_{2},\\ Y(\omega),&\omega\notin A_{2}.\end{cases}$
This is a measurable random closed set with nonempty values, so by measurable selection there exists $y^{+}$ with $y^{+}(\omega)\in G_{+}(\omega)$ almost surely (see [CastaingValadier1977, Molchanov2005]). Then $y^{+}(\omega)\in A$ whenever $Y(\omega)\cap A\neq\emptyset$, and necessarily $y^{+}(\omega)\notin A$ when $Y(\omega)\cap A=\emptyset$. Hence
$\mathbf{P}(y^{+}\in A)=\mathbf{P}(Y\cap A\neq\emptyset),$
so the upper bound in (2.13) is attainable.
For the lower bound, define
$G_{-}(\omega):=\begin{cases}Y(\omega)\setminus A,&\omega\notin A_{1},\ \text{if this set is nonempty},\\ Y(\omega),&\omega\in A_{1}.\end{cases}$
Again $G_{-}$ is a measurable random closed set with nonempty values. A measurable selection $y^{-}$ with $y^{-}(\omega)\in G_{-}(\omega)$ almost surely exists. By construction $y^{-}(\omega)\notin A$ whenever $Y(\omega)\not\subseteq A$, and $y^{-}(\omega)\in A$ if $Y(\omega)\subseteq A$. Thus
$\mathbf{P}(y^{-}\in A)=\mathbf{P}(Y\subseteq A),$
and the lower bound is attained. ∎
We next impose the mean restriction and define, for $\kappa\in[\mathbf{E}y_{L},\mathbf{E}y_{U}]$,
$\operatorname{\mathbf{Sel}}(Y\,|\,\kappa):=\{\,y\in\operatorname{\mathbf{Sel}}(Y):\ \mathbf{E}[y]=\kappa\,\},$
and
$U(\kappa):=\sup_{y\in\operatorname{\mathbf{Sel}}(Y\,|\,\kappa)}\mathbf{P}(y\in A),\quad L(\kappa):=\inf_{y\in\operatorname{\mathbf{Sel}}(Y\,|\,\kappa)}\mathbf{P}(y\in A).$
Introduce the $\omega$-wise extremal points of $A$ inside $Y$ and the associated gaps
$a^{+}(\omega)$ $:=\sup\big(Y(\omega)\cap A\big)\in[-\infty,\infty],$ $a^{-}(\omega)$ $:=\inf\big(Y(\omega)\cap A\big)\in[-\infty,\infty],$ (2.14)
$\Delta^{+}(\omega)$ $:=\big(y_{U}(\omega)-a^{+}(\omega)\big)_{+},$ $\Delta^{-}(\omega)$ $:=\big(a^{-}(\omega)-y_{L}(\omega)\big)_{+},$ (2.15)
with the conventions $\sup\varnothing=$-$\infty$ and $\inf\varnothing=+\infty$.
The next result uses a Lagrangian and an interchange of $\sup$ and expectation in the spirit of [Artstein1974, Molchanov2005]; convex duality arguments are standard, see e.g. [Rockafellar1970].
Theorem 2.13 (Dual characterization under a mean restriction).
Define, for $\lambda\in\mathbb{R}$ and $\omega\in\Omega$,
$\Psi(\lambda,\omega)=\begin{cases}\max\{\lambda y_{U}(\omega),\ 1+\lambda a^{+}(\omega)\},&\lambda\geq 0,\\[2.0pt] \max\{\lambda y_{L}(\omega),\ 1+\lambda a^{-}(\omega)\},&\lambda<0,\end{cases}$
and
$\Phi(\lambda,\omega)=\begin{cases}\min\{\lambda y_{L}(\omega),\ 1+\lambda a^{-}(\omega)\},&\lambda\geq 0,\\[2.0pt] \min\{\lambda y_{U}(\omega),\ 1+\lambda a^{+}(\omega)\},&\lambda<0.\end{cases}$
Then
$U(\kappa)=\inf_{\lambda\in\mathbb{R}}\Big\{\mathbf{E}\big[\Psi(\lambda,\cdot)\big]-\lambda\kappa\Big\},\qquad L(\kappa)=\sup_{\lambda\in\mathbb{R}}\Big\{\mathbf{E}\big[\Phi(\lambda,\cdot)\big]-\lambda\kappa\Big\}.$ (2.16)
Moreover, $U$ is concave and nonincreasing in $\kappa$, while $L$ is convex and nondecreasing.
Proof.
We sketch the argument for $U(\kappa)$; the case of $L(\kappa)$ is analogous.
For each $\lambda\in\mathbb{R}$, consider the Lagrangian functional
$\mathcal{L}(y,\lambda):=\mathbb{E}\big[\mathbf{1}_{A}(y)\big]+\lambda\big(\mathbb{E}[y]-\kappa\big)=\mathbb{E}\big[\mathbf{1}_{A}(y)+\lambda y\big]-\lambda\kappa,\qquad y\in\operatorname{\mathbf{Sel}}(Y).$
For fixed $\lambda$, maximizing $\mathcal{L}(y,\lambda)$ over $y\in\operatorname{\mathbf{Sel}}(Y)$ under the pointwise constraint $y(\omega)\in Y(\omega)$ is equivalent to maximizing
$\mathbb{E}\big[f_{\lambda}(\omega,y(\omega))\big],\qquad f_{\lambda}(\omega,x):=\mathbf{1}_{A}(x)+\lambda x,$
over all selections $y$ of $Y$. By an interchange theorem for random sets (e.g. Artstein [Artstein1974], see also [Molchanov2005]),
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathbb{E}[f_{\lambda}(\omega,y(\omega))]=\mathbb{E}\Big[\sup_{x\in Y(\omega)}f_{\lambda}(\omega,x)\Big].$
For $\lambda\geq 0$,
$\sup_{x\in Y(\omega)}f_{\lambda}(\omega,x)=\max\Big\{\sup_{x\in Y(\omega)\setminus A}\lambda x,\ \sup_{x\in Y(\omega)\cap A}(1+\lambda x)\Big\}=\max\{\lambda y_{U}(\omega),1+\lambda a^{+}(\omega)\},$
with the understanding that the second term is -$\infty$ if $Y(\omega)\cap A=\emptyset$. For $\lambda<0$ one similarly obtains $\max\{\lambda y_{L}(\omega),1+\lambda a^{-}(\omega)\}$. Thus
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathcal{L}(y,\lambda)=\mathbb{E}[\Psi(\lambda,\cdot)]-\lambda\kappa.$
On the other hand, for each $y\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)$, $\mathcal{L}(y,\lambda)=\mathbf{P}(y\in A)$ for all $\lambda$, so
$U(\kappa)=\sup_{y\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)}\mathbf{P}(y\in A)=\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\inf_{\lambda\in\mathbb{R}}\mathcal{L}(y,\lambda).$
By convex duality (the feasible set $\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y)\}$ is the compact interval $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ and the objective is linear in $y$),
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\inf_{\lambda\in\mathbb{R}}\mathcal{L}(y,\lambda)=\inf_{\lambda\in\mathbb{R}}\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathcal{L}(y,\lambda),$
see, e.g., Rockafellar [Rockafellar1970]. Combining this with the calculation above yields the expression for $U(\kappa)$ in (2.16). The expression for $L(\kappa)$ is obtained by applying the same reasoning to the minimization problem for $\mathbf{P}(y\in A)$ subject to $\mathbb{E}[y]=\kappa$. The concavity/convexity and monotonicity properties follow directly from the dual envelopes. ∎
The dual envelopes arise from maximizing/minimizing the Lagrangian $x\mapsto\mathbf{1}\{x\in A\}+\lambda x$ over $x\in Y(\omega)$ and then interchanging expectation and extremum. The corresponding pointwise optimizers have a simple threshold structure.
Theorem 2.14 (Threshold optimizer and calibration).
For each $\lambda\in\mathbb{R}$, let $y_{\lambda}(\omega)\in\arg\max_{x\in Y(\omega)}\big(\mathbf{1}\{x\in A\}+\lambda x\big)$ with measurable tie-breaking on the boundary. Then
$y_{\lambda}(\omega)=\begin{cases}y_{U}(\omega),&\lambda>0,\ Y(\omega)\cap A=\emptyset,\\[4.0pt] y_{U}(\omega)-\min\{\Delta^{+}(\omega),\,1/\lambda\},&\lambda>0,\ Y(\omega)\cap A\neq\emptyset,\\[4.0pt] y_{L}(\omega)+\min\{\Delta^{-}(\omega),\,1/(-\lambda)\},&\lambda<0,\\[4.0pt] \text{any element of }Y(\omega),&\lambda=0.\end{cases}$ (2.17)
The map $\lambda\mapsto\mathbf{E}[y_{\lambda}]$ is continuous and nondecreasing, with $\lim_{\lambda\to-\infty}\mathbf{E}[y_{\lambda}]=\mathbf{E}y_{L}$ and $\lim_{\lambda\to+\infty}\mathbf{E}[y_{\lambda}]=\mathbf{E}y_{U}$. Hence for each $\kappa\in[\mathbf{E}y_{L},\mathbf{E}y_{U}]$ there exists $\lambda^{\star}$ such that $\mathbf{E}[y_{\lambda^{\star}}]=\kappa$, and any such $y_{\lambda^{\star}}$ attains $U(\kappa)$.
Proof.
Fix $\lambda>0$. On each $\omega$, the function $x\mapsto\mathbf{1}_{A}(x)+\lambda x$ is affine with slope $\lambda>0$ outside $A$, and affine with the same slope plus a unit jump on $A$. If $Y(\omega)\cap A=\emptyset$, the maximizer is clearly $y_{U}(\omega)$, giving the first line of (2.17). If $Y(\omega)\cap A\neq\emptyset$, denote
$a^{+}(\omega)=\sup(Y(\omega)\cap A),\qquad\Delta^{+}(\omega)=y_{U}(\omega)-a^{+}(\omega)\geq 0.$
On $Y(\omega)\cap A$, the maximum of $1+\lambda x$ is attained at $x=a^{+}(\omega)$, giving value $1+\lambda a^{+}(\omega)$. On $Y(\omega)$, the maximum of $\lambda x$ is $\lambda y_{U}(\omega)$. Thus the maximizer is $y_{U}(\omega)$ if $\lambda y_{U}(\omega)\geq 1+\lambda a^{+}(\omega)$, i.e. $\lambda\Delta^{+}(\omega)\geq 1$; otherwise it is $a^{+}(\omega)$. This yields
$y_{\lambda}(\omega)=\begin{cases}y_{U}(\omega),&\lambda\Delta^{+}(\omega)\geq 1,\\ a^{+}(\omega),&\lambda\Delta^{+}(\omega)<1,\end{cases}$
which is equivalent, for $Y(\omega)\cap A\neq\emptyset$, to
$y_{\lambda}(\omega)=y_{U}(\omega)-\min\{\Delta^{+}(\omega),\,1/\lambda\}$
when $\lambda>0$. The case $\lambda<0$ is analogous, with roles of $y_{U}$ and $y_{L}$ reversed and $a^{-}(\omega)=\inf(Y(\omega)\cap A)$.
As $\lambda\to+\infty$, $1/\lambda\to 0$ and hence $\min\{\Delta^{+},1/\lambda\}\to 0$ almost surely on $\{Y\cap A\neq\emptyset\}$, while $y_{\lambda}(\omega)=y_{U}(\omega)$ when $Y(\omega)\cap A=\emptyset$, so $y_{\lambda}\to y_{U}$ and $\mathbb{E}[y_{\lambda}]\to\mathbb{E}[y_{U}]$. As $\lambda\to$-$\infty$, $y_{\lambda}\to y_{L}$ and $\mathbb{E}[y_{\lambda}]\to\mathbb{E}[y_{L}]$. Monotonicity and continuity of $\lambda\mapsto\mathbb{E}[y_{\lambda}]$ follow from the monotone dependence of the thresholds on $\lambda$ and dominated convergence. Existence of a $\lambda^{\star}$ with $\mathbb{E}[y_{\lambda^{\star}}]=\kappa$ for any $\kappa\in[\mathbb{E}y_{L},\mathbb{E}y_{U}]$ then follows by the intermediate value theorem.
Finally, for such a $\lambda^{\star}$, the dual expression in Theorem 2.13 yields
$U(\kappa)\geq\mathbf{P}(y_{\lambda^{\star}}\in A)=\mathbb{E}\big[\mathbf{1}_{A}(y_{\lambda^{\star}})\big],$
while the Lagrangian argument shows that equality holds at the calibrated $\lambda^{\star}$. ∎
The next result identifies the optimal probability in closed form once the calibration threshold is determined.
Corollary 2.15 (Identification formula for $\mathbf{P}(y^{*}\in A)$).
Let $y^{\star}:=y_{\lambda^{\star}}$ be an optimizer from Theorem˜2.14, and set $t^{\star}:=1/\lambda^{\star}$ if $\lambda^{\star}>0$ and $s^{\star}:=1/(-\lambda^{\star})$ if $\lambda^{\star}<0$. Then
[…]
and the thresholds are pinned down by the mean-calibration equations
$\mathbf{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big]=\kappa\quad(\lambda^{\star}>0),\qquad\mathbf{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big]=\kappa\quad(\lambda^{\star}<0).$ (2.19)
Proof.
For $\lambda^{\star}>0$, the optimizer in Theorem 2.14 has the form
$y^{\star}(\omega)=\begin{cases}y_{U}(\omega),&Y(\omega)\cap A=\emptyset,\\[2.0pt] y_{U}(\omega)-\min\{\Delta^{+}(\omega),t^{\star}\},&Y(\omega)\cap A\neq\emptyset,\end{cases}$
with $t^{\star}=1/\lambda^{\star}$. On $\{Y\cap A\neq\emptyset\}$, when $\Delta^{+}(\omega)<t^{\star}$ we have $y^{\star}(\omega)=a^{+}(\omega)\in A$; when $\Delta^{+}(\omega)>t^{\star}$ we have $y^{\star}(\omega)=y_{U}(\omega)\notin A$ (except possibly on the boundary of $A$). On the level set $\{\Delta^{+}=t^{\star}\}$ both choices yield the same value of the Lagrangian, so tie-breaking can be used to match the mean constraint while keeping the probability as in (2.18). The mean calibration (2.19) is exactly the equation $\mathbb{E}[y^{\star}]=\kappa$ restricted to the case $\lambda^{\star}>0$. The case $\lambda^{\star}<0$ is analogous, using the expression for $y_{\lambda}$ in terms of $y_{L}$ and $\Delta^{-}$. When $\lambda^{\star}=0$, the mean constraint is nonbinding in the Lagrangian and the optimal strategy is simply to choose a selection that hits $A$ whenever possible, giving probability $\mathbf{P}(Y\cap A\neq\emptyset)$. ∎
Finally, we record qualitative properties and a convenient specialization.
Proposition 2.16 (Extremes and global consistency).
At the mean extremes,
$U(\mathbf{E}y_{U})=L(\mathbf{E}y_{U})=\mathbf{P}\big(y_{U}\in A\big),\qquad U(\mathbf{E}y_{L})=L(\mathbf{E}y_{L})=\mathbf{P}\big(y_{L}\in A\big),$
and for all $\kappa\in[\mathbf{E}y_{L},\mathbf{E}y_{U}]$ the global random-set bounds are preserved:
$\mathbf{P}\!\big(Y\subseteq A\big)\ \leq\ L(\kappa)\ \leq\ U(\kappa)\ \leq\ \mathbf{P}\!\big(Y\cap A\neq\varnothing\big).$
Proof.
When $\kappa=\mathbb{E}y_{U}$, the mean constraint forces $y=y_{U}$ almost surely, so $U(\kappa)=L(\kappa)=\mathbf{P}(y_{U}\in A)$; similarly for $\kappa=\mathbb{E}y_{L}$. The global bounds follow because the mean restriction can only shrink the feasible set of selections, and Proposition 2.12 applies to any $y\in\operatorname{\mathbf{Sel}}(Y\mid\kappa)$. ∎
Remark 2.17 (Interval targets).
When $A=[a,b]$,
$\Delta^{+}(\omega)=\begin{cases}\big(y_{U}(\omega)-b\big)_{+},&y_{L}(\omega)\leq b,\\ +\infty,&y_{L}(\omega)>b,\end{cases}\qquad\Delta^{-}(\omega)=\begin{cases}\big(a-y_{L}(\omega)\big)_{+},&y_{U}(\omega)\geq a,\\ +\infty,&y_{U}(\omega)<a,\end{cases}$
so the threshold rule (2.17) switches at a deterministic distance from the right edge $b$ in the high-mean regime and from the left edge $a$ in the low-mean regime, with $t^{\star}$ or $s^{\star}$ determined by (2.19).
2.8. Further extensions: higher moments and additional quantiles
We now indicate how the selection-based approach developed above extends to higher moments and to quantile restrictions other than the median. We keep the standing assumptions of Subsection 1.3 and of Section 2: $Y=[y_{L},y_{U}]$ is a random interval with $\mathbf{P}(y_{L}\leq y_{U})=1$ and $\mathbb{E}(|y_{L}|+|y_{U}|)<\infty$ on a non-atomic probability space $(\Omega,\mathcal{A},\mathbf{P})$.
Higher moments
We first consider restrictions on the $r$th moment of selections for some fixed $r\in\mathbb{R}$.
Definition 2.18 (Moment-restricted selection sets).
Let $r\in\mathbb{R}$ and $\mu_{r}\in\mathbb{R}$. The $r$th-moment restricted selection set is
$\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r}):=\bigl\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}|y|^{r}<\infty,\ \mathbb{E}[y^{r}]=\mu_{r}\bigr\}.$
When no confusion arises we write $\operatorname{\mathbf{Sel}}_{r}(\mu_{r})$ for $\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})$.
We begin with a simple non-emptiness result in the case where the pointwise power map $x\mapsto x^{r}$ preserves the interval structure of $Y$.
Proposition 2.19 (Non-emptiness under power transformations).
Suppose either
  1. (i)
    $r$ is an odd integer, or
  2. (ii)
    $r>0$ and $y_{L}\geq 0$ almost surely.
Define the random set
$Y^{(r)}(\omega):=\{x^{r}:x\in Y(\omega)\},\qquad\omega\in\Omega.$
Then $Y^{(r)}$ is a random interval, and
$\mathbb{E}\bigl[Y^{(r)}\bigr]=\bigl[\mathbb{E}(y_{L}^{r}),\mathbb{E}(y_{U}^{r})\bigr]=\overline{\bigl\{\mathbb{E}[y^{r}]:y\in\operatorname{\mathbf{Sel}}(Y)\bigr\}}.$
In particular, $\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})\neq\emptyset$ if and only if $\mu_{r}\in[\mathbb{E}(y_{L}^{r}),\mathbb{E}(y_{U}^{r})]$.
Proof.
Under (i) the map $x\mapsto x^{r}$ is strictly increasing on $\mathbb{R}$, so $Y^{(r)}(\omega)=[y_{L}(\omega)^{r},y_{U}(\omega)^{r}]$ for all $\omega$. Under (ii) the map is increasing on $0,\infty)$ and $Y(\omega)\subseteq[0,\infty)$ almost surely, so again $Y^{(r)}(\omega)=[y_{L}(\omega)^{r},y_{U}(\omega)^{r}]$. In both cases $Y^{(r)}$ is a random interval with endpoints $L_{r}:=y_{L}^{r}$ and $U_{r}:=y_{U}^{r}$. The Aumann expectation identity (1.4) applied to $Y^{(r)}$ yields
$\mathbb{E}\bigl[Y^{(r)}\bigr]=\bigl[\mathbb{E}(y_{L}^{r}),\mathbb{E}(y_{U}^{r})\bigr].$
If $y\in\operatorname{\mathbf{Sel}}(Y)$, then $y^{r}$ is a measurable selection of $Y^{(r)}$, so $\mathbb{E}[y^{r}]\in\mathbb{E}[Y^{(r)}]$. Conversely, for any $m\in\mathbb{E}[Y^{(r)}]$ there exists $z\in\operatorname{\mathbf{Sel}}(Y^{(r)})$ with $\mathbb{E}[z]=m$. Because $x\mapsto x^{1/r}$ is strictly increasing on the relevant domain (all of $\mathbb{R}$ in case (i), $0,\infty)$ in case (ii)), the random variable $y:=z^{1/r}$ is a measurable selection of $Y$ satisfying $y^{r}=z$ and hence $\mathbb{E}[y^{r}]=\mathbb{E}[z]=m$. Thus the closure of $\{\mathbb{E}[y^{r}]:y\in\operatorname{\mathbf{Sel}}(Y)\}$ equals $\mathbb{E}[Y^{(r)}]$, and the final statement follows from Definition 2.18. ∎
We now describe the induced range of means under an $r$th-moment restriction. Define, for $\mu_{r}\in[\mathbb{E}(y_{L}^{r}),\mathbb{E}(y_{U}^{r})]$,
$\underline{m}(r,\mu_{r}):=\inf_{y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})}\mathbb{E}[y],\qquad\overline{m}(r,\mu_{r}):=\sup_{y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})}\mathbb{E}[y],$
with the convention that $\underline{m}(r,\mu_{r})=+\infty$, $\overline{m}(r,\mu_{r})=$-$\infty$ if $\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})=\emptyset$.
Theorem 2.20 (Dual representation under an $r$th-moment restriction).
Let the assumptions of Proposition 2.19 hold and let $\mu_{r}\in[\mathbb{E}(y_{L}^{r}),\mathbb{E}(y_{U}^{r})]$. For $\lambda\in\mathbb{R}$ and $\omega\in\Omega$ define
$\Psi_{r}(\lambda,\omega):=\sup_{x\in Y(\omega)}\bigl(x+\lambda x^{r}\bigr),\qquad\Phi_{r}(\lambda,\omega):=\inf_{x\in Y(\omega)}\bigl(x+\lambda x^{r}\bigr).$
Then
$\overline{m}(r,\mu_{r})=\inf_{\lambda\in\mathbb{R}}\Bigl\{\mathbb{E}\bigl[\Psi_{r}(\lambda,\cdot)\bigr]-\lambda\mu_{r}\Bigr\},\qquad\underline{m}(r,\mu_{r})=\sup_{\lambda\in\mathbb{R}}\Bigl\{\mathbb{E}\bigl[\Phi_{r}(\lambda,\cdot)\bigr]-\lambda\mu_{r}\Bigr\}.$ (2.20)
In particular, the $r$th-moment restricted mean set $\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})\}$ is a nonempty compact interval $\underline{m}(r,\mu_{r}),\overline{m}(r,\mu_{r})]\subseteq[\mathbb{E}y_{L},\mathbb{E}y_{U}]$.
Proof.
We give the argument for $\overline{m}(r,\mu_{r})$; the lower endpoint is handled analogously.
Consider the optimization problem
$\sup\bigl\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}(Y),\ \mathbb{E}[y^{r}]=\mu_{r}\bigr\}.$
Introduce a Lagrange multiplier $\lambda\in\mathbb{R}$ for the moment constraint and define the Lagrangian
$\mathcal{L}(y,\lambda):=\mathbb{E}[y]+\lambda\bigl(\mathbb{E}[y^{r}]-\mu_{r}\bigr)=\mathbb{E}\bigl[y+\lambda y^{r}\bigr]-\lambda\mu_{r},\qquad y\in\operatorname{\mathbf{Sel}}(Y).$
For fixed $\lambda$, maximizing $\mathcal{L}(\cdot,\lambda)$ over $\operatorname{\mathbf{Sel}}(Y)$ subject to $y(\omega)\in Y(\omega)$ is equivalent to maximizing
$\mathbb{E}\bigl[f_{\lambda}(\omega,y(\omega))\bigr],\qquad f_{\lambda}(\omega,x):=x+\lambda x^{r},$
over selections $y$ of $Y$. By an Artstein-type interchange theorem for random closed sets (see [Artstein1974, Molchanov2005]),
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathbb{E}\bigl[f_{\lambda}(\omega,y(\omega))\bigr]=\mathbb{E}\Bigl[\sup_{x\in Y(\omega)}f_{\lambda}(\omega,x)\Bigr]=\mathbb{E}\bigl[\Psi_{r}(\lambda,\cdot)\bigr].$
Therefore
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathcal{L}(y,\lambda)=\mathbb{E}\bigl[\Psi_{r}(\lambda,\cdot)\bigr]-\lambda\mu_{r}.$
On the other hand, for any $y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})$ one has $\mathcal{L}(y,\lambda)=\mathbb{E}[y]$ for all $\lambda$, so
$\overline{m}(r,\mu_{r})=\sup_{y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})}\mathbb{E}[y]=\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\inf_{\lambda\in\mathbb{R}}\mathcal{L}(y,\lambda).$
The feasible set of pairs
$\bigl(\mathbb{E}[y],\mathbb{E}[y^{r}]\bigr),\qquad y\in\operatorname{\mathbf{Sel}}(Y),$
is the compact convex subset of $\mathbb{R}^{2}$ given by the Aumann expectation of the random closed set
$\Xi(\omega):=\{(x,x^{r}):x\in Y(\omega)\},$
cf. [Artstein1974, Molchanov2005]. Standard convex duality arguments (see, e.g., [Rockafellar1970]) yield
$\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\inf_{\lambda\in\mathbb{R}}\mathcal{L}(y,\lambda)=\inf_{\lambda\in\mathbb{R}}\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathcal{L}(y,\lambda),$
so
$\overline{m}(r,\mu_{r})=\inf_{\lambda\in\mathbb{R}}\sup_{y\in\operatorname{\mathbf{Sel}}(Y)}\mathcal{L}(y,\lambda)=\inf_{\lambda\in\mathbb{R}}\Bigl\{\mathbb{E}\bigl[\Psi_{r}(\lambda,\cdot)\bigr]-\lambda\mu_{r}\Bigr\},$
which is the first identity in (2.20). The expression for $\underline{m}(r,\mu_{r})$ follows analogously, by applying the same argument to the minimization problem for $\mathbb{E}[y]$ with $\mathbb{E}[y^{r}]=\mu_{r}$.
Finally, because the set of feasible pairs $(\mathbb{E}[y],\mathbb{E}[y^{r}])$ is compact and convex, the intersection with the affine constraint $\mathbb{E}[y^{r}]=\mu_{r}$ is either empty or a compact interval in the first coordinate. Under the non-emptiness condition of Proposition 2.19, the attainable mean set $\{\mathbb{E}[y]:y\in\operatorname{\mathbf{Sel}}_{r}(Y\mid\mu_{r})\}$ is a nonempty compact interval contained in $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ by (1.4). ∎
Remark 2.21.
For $r=1$ and $\mu_{1}=\kappa$ we have $\operatorname{\mathbf{Sel}}_{1}(Y\mid\mu_{1})=\operatorname{\mathbf{Sel}}(Y\mid\kappa)$ and $\underline{m}(1,\mu_{1})=\overline{m}(1,\mu_{1})=\kappa$, recovering the trivial point identification of the mean. For $r\neq 1$, the interval $\underline{m}(r,\mu_{r}),\overline{m}(r,\mu_{r})]$ describes the shrinkage of the benchmark range $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ induced by the $r$th-moment restriction.
Additional quantile restrictions
We next consider quantile restrictions away from the median. Recall from Proposition 2.4 that for each $\alpha\in(0,1)$ and each $q_{\alpha}$ in the feasible range
$q_{\alpha}\in\bigl[T_{Y}^{-1}(\alpha),\,C_{Y}^{-1}(\alpha)\bigr],$
there exists a selection $y\in\operatorname{\mathbf{Sel}}(Y)$ such that $F_{y}^{-1}(\alpha)=q_{\alpha}$.
Definition 2.22 (Quantile-restricted selection sets).
Fix $\alpha\in(0,1)$ and $q_{\alpha}\in\mathbb{R}$. The $(\alpha,q_{\alpha})$-quantile restricted selection set is
$\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha}):=\bigl\{y\in\operatorname{\mathbf{Sel}}(Y):F_{y}^{-1}(\alpha)=q_{\alpha}\bigr\},$
where $F_{y}^{-1}$ denotes the left-continuous quantile function of $y$.
We are interested in the induced range of means
$\Theta_{E}(\alpha,q_{\alpha}):=\bigl\{\mathbb{E}[y]:\,y\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})\bigr\}.$
We first record existence and basic bounds, and then show that $\Theta_{E}(\alpha,q_{\alpha})$ is a convex subset of $\mathbb{E}y_{L},\mathbb{E}y_{U}]$.
Proposition 2.23 (Non-emptiness and basic bounds).
Let $\alpha\in(0,1)$ and assume
$q_{\alpha}\in\bigl[T_{Y}^{-1}(\alpha),\,C_{Y}^{-1}(\alpha)\bigr].$
Then $\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})\neq\emptyset$, and every $y\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ satisfies
$\mathbb{E}[y]\ \in\ \bigl[\mathbb{E}(y_{L}),\,\mathbb{E}(y_{U})\bigr].$
In particular,
$\emptyset\neq\Theta_{E}(\alpha,q_{\alpha})\ \subseteq\ [\mathbb{E}y_{L},\mathbb{E}y_{U}].$
Proof.
By Proposition 2.4, if $q_{\alpha}\in[T_{Y}^{-1}(\alpha),C_{Y}^{-1}(\alpha)]$ then there exists a selection $y\in\operatorname{\mathbf{Sel}}(Y)$ such that $F_{y}^{-1}(\alpha)=q_{\alpha}$. This shows $\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})\neq\emptyset$.
For any $y\in\operatorname{\mathbf{Sel}}(Y)$ we have $y_{L}\leq y\leq y_{U}$ almost surely, hence by monotonicity of expectation,
$\mathbb{E}(y_{L})\ \leq\ \mathbb{E}[y]\ \leq\ \mathbb{E}(y_{U}).$
This applies in particular to $y\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$, so
$\Theta_{E}(\alpha,q_{\alpha})\subseteq[\mathbb{E}y_{L},\mathbb{E}y_{U}].$
Together with non-emptiness this yields the claim. ∎
The quantile restriction is non-linear in $y$ pointwise, so $y^{(1)},y^{(2)}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ does not imply that their convex combination $\theta y^{(1)}+(1-\theta)y^{(2)}$ has the same $\alpha$-quantile. Convexity of $\Theta_{E}(\alpha,q_{\alpha})$ therefore has to be established indirectly, at the level of distributions and by allowing exogenous randomization on a product extension of the underlying probability space.
We first note that the class of distributions with a fixed $\alpha$-quantile is convex.
Lemma 2.24 (Convexity of the fixed-quantile constraint in law).
Let $\alpha\in(0,1)$ and $q\in\mathbb{R}$. Let $F_{1},F_{2}$ be distribution functions on $\mathbb{R}$ such that
$F_{i}^{-1}(\alpha)\ =\ q,\qquad i=1,2,$
where $F_{i}^{-1}(\alpha):=\inf\{t\in\mathbb{R}:F_{i}(t)\geq\alpha\}$ is the left-continuous $\alpha$-quantile. For $\theta\in[0,1]$ define
$F_{\theta}:=\theta F_{1}+(1-\theta)F_{2}.$
Then $F_{\theta}^{-1}(\alpha)=q$.
Proof.
Write $q_{\alpha}(F)$ for the $\alpha$-quantile of a distribution function $F$. The condition $q_{\alpha}(F_{i})=q$ means
$F_{i}(t)<\alpha\quad\text{for all }t<q,\qquad F_{i}(q)\geq\alpha,\qquad i=1,2.$
Fix $t<q$. Then $F_{i}(t)<\alpha$ for $i=1,2$, so
$F_{\theta}(t)=\theta F_{1}(t)+(1-\theta)F_{2}(t)<\theta\alpha+(1-\theta)\alpha=\alpha.$
At $t=q$,
$F_{\theta}(q)=\theta F_{1}(q)+(1-\theta)F_{2}(q)\geq\theta\alpha+(1-\theta)\alpha=\alpha.$
Thus $F_{\theta}(t)<\alpha$ for all $t<q$ and $F_{\theta}(q)\geq\alpha$. By the definition of the left-continuous quantile,
$F_{\theta}^{-1}(\alpha)=\inf\{t\in\mathbb{R}:F_{\theta}(t)\geq\alpha\}=q.$
We now use this to show that $\Theta_{E}(\alpha,q_{\alpha})$ is convex. The proof formally works on a product extension that introduces independent randomization, but because our identification analysis only depends on the joint law of $(Y,y)$, the attainable range of $\mathbb{E}[y]$ is unaffected by such an enlargement.
Proposition 2.25 (Convexity of the mean range).
Let $\alpha\in(0,1)$ and $q_{\alpha}$ be as in Proposition 2.23. Then $\Theta_{E}(\alpha,q_{\alpha})$ is a convex subset of $\mathbb{R}$, hence an interval (possibly degenerate) contained in $\mathbb{E}y_{L},\mathbb{E}y_{U}]$. Equivalently, for any $y^{(1)},y^{(2)}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ and any $\theta\in[0,1]$, there exists $\tilde{y}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ such that
$\mathbb{E}[\tilde{y}]=\theta\,\mathbb{E}[y^{(1)}]+(1-\theta)\,\mathbb{E}[y^{(2)}].$
Proof.
Let $y^{(1)},y^{(2)}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ and $\theta\in[0,1]$. Denote by $F_{i}$ the distribution function of $y^{(i)}$, $i=1,2$. By definition of $\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$,
$F_{i}^{-1}(\alpha)=q_{\alpha},\qquad i=1,2.$
Step 1: product extension and randomization. Consider the product probability space
$(\bar{\Omega},\bar{\mathcal{A}},\bar{\mathbf{P}}):=(\Omega\times[0,1],\ \mathcal{A}\otimes\mathcal{B}([0,1]),\ \mathbf{P}\otimes\lambda),$
where $\lambda$ denotes Lebesgue measure on $[0,1]$. Let $U$ be the canonical $\mathrm{Unif}[0,1]$ random variable on $[0,1]$, independent of all random objects defined on $(\Omega,\mathcal{A},\mathbf{P})$. Extend the random interval $Y$ to $\bar{\Omega}$ by
$Y(\omega,u):=Y(\omega),\qquad(\omega,u)\in\bar{\Omega}.$
Each $y^{(i)}$ can be viewed as a random variable on $\bar{\Omega}$ by ignoring the second coordinate:
$y^{(i)}(\omega,u):=y^{(i)}(\omega),\qquad i=1,2.$
Then $y^{(i)}(\omega,u)\in Y(\omega,u)$ for all $(\omega,u)$, so $y^{(i)}\in\operatorname{\mathbf{Sel}}(Y)$ on the extended space as well.
Step 2: mixture selection. Define a randomized selection $\tilde{y}:\bar{\Omega}\to\mathbb{R}$ by
$\tilde{y}(\omega,u):=\begin{cases}y^{(1)}(\omega),&u\leq\theta,\\[4.0pt] y^{(2)}(\omega),&u>\theta.\end{cases}$
For each $(\omega,u)$, $\tilde{y}(\omega,u)$ equals either $y^{(1)}(\omega)$ or $y^{(2)}(\omega)$, so $\tilde{y}(\omega,u)\in Y(\omega,u)$ and $\tilde{y}$ is a measurable selection of $Y$ on $\bar{\Omega}$.
By independence of $U$ and $(y^{(1)},y^{(2)})$, the distribution of $\tilde{y}$ under $\bar{\mathbf{P}}$ is the mixture
$\text{law}(\tilde{y})=\theta\,\text{law}(y^{(1)})+(1-\theta)\,\text{law}(y^{(2)}),$
so its distribution function is
$F_{\theta}(t):=\theta F_{1}(t)+(1-\theta)F_{2}(t),\qquad t\in\mathbb{R}.$
Step 3: quantile and mean of the mixture. Since $F_{i}^{-1}(\alpha)=q_{\alpha}$ for $i=1,2$, Lemma 2.24 applied with $q=q_{\alpha}$ yields
$F_{\theta}^{-1}(\alpha)=q_{\alpha}.$
Thus $\tilde{y}$ satisfies the same $(\alpha,q_{\alpha})$-quantile restriction: $F_{\tilde{y}}^{-1}(\alpha)=q_{\alpha}$, i.e. $\tilde{y}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ when viewed on the extended space. Its mean under $\bar{\mathbf{P}}$ is
$\mathbb{E}_{\bar{\mathbf{P}}}[\tilde{y}]=\theta\,\mathbb{E}_{\mathbf{P}}[y^{(1)}]+(1-\theta)\,\mathbb{E}_{\mathbf{P}}[y^{(2)}],$
because $U$ is independent and selects $y^{(1)}$ with probability $\theta$ and $y^{(2)}$ with probability $1-\theta$.
[…]
$\hat{y}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha}),\qquad\mathbb{E}[\hat{y}]=\mathbb{E}_{\bar{\mathbf{P}}}[\tilde{y}]=\theta\,\mathbb{E}[y^{(1)}]+(1-\theta)\,\mathbb{E}[y^{(2)}].$
Since $y^{(1)},y^{(2)}$ and $\theta\in[0,1]$ were arbitrary, this shows that $\Theta_{E}(\alpha,q_{\alpha})$ is convex. Together with Proposition 2.23, which ensures that $\Theta_{E}(\alpha,q_{\alpha})$ is nonempty and contained in $\mathbb{E}y_{L},\mathbb{E}y_{U}]$, we conclude that $\Theta_{E}(\alpha,q_{\alpha})$ is an interval (possibly a singleton) contained in $\mathbb{E}y_{L},\mathbb{E}y_{U}]$. ∎
It is convenient to summarize the induced mean range under a single quantile restriction by its lower and upper endpoints.
Definition 2.26 (Quantile-identified mean interval).
For $\alpha\in(0,1)$ and $q_{\alpha}$ as above, define
$\underline{m}(\alpha,q_{\alpha}):=\inf\Theta_{E}(\alpha,q_{\alpha}),\qquad\overline{m}(\alpha,q_{\alpha}):=\sup\Theta_{E}(\alpha,q_{\alpha}),$
with the convention that $\underline{m}(\alpha,q_{\alpha})=+\infty$ and $\overline{m}(\alpha,q_{\alpha})=$-$\infty$ if $\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})=\emptyset$. Whenever $\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})\neq\emptyset$, the closed interval
$\bigl[\underline{m}(\alpha,q_{\alpha}),\,\overline{m}(\alpha,q_{\alpha})\bigr]\subseteq[\mathbb{E}y_{L},\mathbb{E}y_{U}]$
will be called the $(\alpha,q_{\alpha})$-quantile identified mean interval.
By Proposition 2.23 and Proposition 2.25, $\Theta_{E}(\alpha,q_{\alpha})$ is a nonempty convex subset of $\mathbb{E}y_{L},\mathbb{E}y_{U}]$, so its infimum and supremum are finite and satisfy $\underline{m}(\alpha,q_{\alpha})\leq\overline{m}(\alpha,q_{\alpha})$. The endpoints need not be attained by a selection in general, but for any $\varepsilon>0$ there exist $y_{\varepsilon}^{-},y_{\varepsilon}^{+}\in\operatorname{\mathbf{Sel}}_{\alpha}(Y\mid q_{\alpha})$ such that
$\mathbb{E}[y_{\varepsilon}^{-}]\leq\underline{m}(\alpha,q_{\alpha})+\varepsilon,\qquad\mathbb{E}[y_{\varepsilon}^{+}]\geq\overline{m}(\alpha,q_{\alpha})-\varepsilon.$
Remark 2.27.
For $\alpha=\tfrac{1}{2}$ and a singleton median $q_{1/2}=m$, the set $\operatorname{\mathbf{Sel}}_{1/2}(Y\mid m)$ coincides with $\operatorname{\mathbf{Sel}}(Y\mid m)$ from Subsection 2.5, and the mean range $\Theta_{E}(1/2,m)$ is exactly the interval $\mathbb{E}_{\min}(m),\mathbb{E}_{\max}(m)]$ characterized in Proposition 2.7. For general $\alpha\in(0,1)$, Propositions 2.23 and 2.25 show that a single quantile restriction always leads to a well-defined convex identified set of means contained in the Aumann range $\mathbb{E}y_{L},\mathbb{E}y_{U}]$. Extending the explicit quantile-integral formulas of Proposition 2.7 to arbitrary levels $\alpha\neq\tfrac{1}{2}$ would require a more refined analysis of the tails below and above $q_{\alpha}$ and additional regularity assumptions; we do not pursue this here.
Credit authorship contribution statement:
Arie Beresteanu: Writing - review & editing, Writing - original draft, Methodology, Investigation, Conceptualization.
Behrooz Moosavi Ramezanzadeh: Writing - review & editing, Writing - original draft, Methodology, Investigation, Conceptualization.